### 30 Mar 2026 v1

# Incompressible Euler equations in 3D bounded domains in a critical space

Abstract.
We consider the 3D incompressible Euler equations in bounded domains $\Omega$ with smooth boundary $\partial\Omega$. Based on [IMT-2019], we define the Besov space $B^{s}_{p,q}(A)$ by means of the Stokes operator $A$ with the Neumann boundary condition on $\partial\Omega$, and prove unique local existence theorem of strong solution for the initial data in the critical Besov space $B^{\frac{5}{2}}_{2,1}(A)$. Our proof relies on the method of vanishing viscosity. The commutator estimate plays an essential role for derivation of energy bounds which hold uniformly with respect to viscosity constants.
Key words and phrases:
Euler equations; bounded domain; critical Besov space; spectral resolution; Neumann boundary condition; vanishing viscosity
1. Introduction
0002010 Mathematics Subject Classification: 35Q35; 35Q86
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$ with the smooth boundary $\partial\Omega$. In $\Omega$ we consider the following initial-boundary value problem of the Euler equations for the incomressible perfect fluid;
$\begin{cases}\partial_{t}u+(u\cdot\nabla)u+\nabla\pi=0&\mbox{in $\Omega\times(0,T)$},\\ {\rm div\,}u=0&\mbox{in $\Omega\times(0,T)$},\\ u\cdot\nu=0&\mbox{on $\partial\Omega\times(0,T)$},\\ u(x,0)=u_{0}(x)&\mbox{in $\Omega$},\end{cases}$ (1.1)
where $u=u(x,t)=(u_{1}(x,t),u_{2}(x,t),u_{3}(x,t))$ and $\pi=\pi(x,t)$ denote the unknown velocity vector and pressure of the fluid at the point $x=(x_{1},x_{2},x_{3})\in\Omega$ and the time $t\in(0,T)$, respectively, while $u_{0}=u_{0}(x)=(u_{0,1}(x),u_{0,2}(x),u_{0,3}(x))$ is the given initial velocity vector. Here $\nu=(\nu_{1},\nu_{2},\nu_{3})$ denotes the unit outer normal to $\partial\Omega$.
The purpose of this paper is to prove local existence theorem of classical solution $u$ to (1.1) for arbitrary $u_{0}$ in a critical Besov space on $\Omega$. Temam [Te-1975] proved local existence of strong solutions of (1.1) with $u_{0}\in W^{m,p}(\Omega)$ for $1<p<\infty$ and an integer $m>1+3/p$. In the higher dimensional case, the corresponding result was obtained by Bourguignon-Brezis [BoBr-1974]. Kato-Lai [KaLa-1984] handled the case $\Omega\subset\mathbb{R}^{n}$ and established an existence theorem of local solutions to abstract evolution equations including (1.1) with $u_{0}\in H^{m}(\Omega)$ for an integer $m\geq 2+[n/2]$. On the other hand, in the case when $\Omega=\mathbb{R}^{n}$, there are a number of literatures in the frame work of Besov spaces $B^{s}_{p,q}(\mathbb{R}^{n})$ and Triebel-Lisorkin spaces $F^{s}_{p,q}(\mathbb{R}^{n})$. Vishik [Vi-1999] constructed a local solution with the initial vorticity $\operatorname{rot}u_{0}\in B^{s}_{p,q}(\mathbb{R}^{n})$ for $1<p,q<\infty$ and $s>n/p$. Chae [Ch-2002] obtained the corresponding result to the initial data $u_{0}\in F^{s}_{p,q}(\mathbb{R}^{n})$ for $1<p,q<\infty$ and $s>1+n/p$. Pak-Park [PaPa-2004] has been successful to obtain the local solution in such a critical Besov space as $B^{1}_{\infty,1}(\mathbb{R}^{n})$.
In the present paper, we prove the local well-posedness in the critical Besov space $B^{\frac{5}{2}}_{2,1}(\Omega)$. Based on [IwKo-preprint] and [IMT-2019], we first define the Besov space $B^{s}_{p,q}(A)$ in terms of the Stokes operator $A$ in $L^{2}_{\sigma}(\Omega)$ with the Neumann boundary condition on $\partial\Omega$. Our method relies on the invisid limit ${\lim_{\varepsilon\to+0}u_{\varepsilon}}$ of a family $\{u_{\varepsilon}\}_{\varepsilon>0}$ of solutions to the following Navier-Stokes equations;
$\begin{cases}\partial_{t}u_{\varepsilon}-\varepsilon\Delta u_{\varepsilon}+(u_{\varepsilon}\cdot\nabla)u_{\varepsilon}+\nabla\pi_{\varepsilon}=0&\mbox{in $\Omega\times(0,T)$},\\ {\rm div\,}u_{\varepsilon}=0&\mbox{in $\Omega\times(0,T)$},\\ u_{\varepsilon}\cdot\nu=0,\,\,\,\operatorname{rot}u_{\varepsilon}\times\nu=0&\mbox{on $\partial\Omega\times(0,T)$},\\ u_{\varepsilon}(x,0)=u_{0}(x)&\mbox{in $\Omega$}.\end{cases}$ (1.2)
It is essential to establish uniform bounds of $\{u_{\varepsilon}\}_{\varepsilon>0}$ in the Chemin-Lernar space $\tilde{L}^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$. The commutator type estimate such as Kato-Ponce [KaPo-1988] plays a crucial role for such bounds. Since we make use of the Stokes operator $A$ with the Neumann boundary condition like (1.2), the boundary layer does not appear in the invisid limit.
To state our result, let us recall the Stokes operator $A_{p}$ on $L^{p}_{\sigma}(\Omega)$ for $1<p<\infty$. We denote by $C^{\infty}_{0,\sigma}(\Omega)$ the set of $C^{\infty}$-solenoidal vector functions with compact support in $\Omega$. $L^{p}_{\sigma}(\Omega)$ is the closure of $C^{\infty}_{0,\sigma}(\Omega)$ with respect to the $L^{p}$-norm $\|\cdot\|_{L^{p}}$ on $\Omega$. Then the Helmholtz projection $\mathbb{P}_{p}$ from $L^{p}(\Omega)$ onto $L^{p}_{\sigma}(\Omega)$ is well-defined. We define $A_{p}$ by
[…]
Concerning the kernel $\mbox{Ker}A_{p}$ of $A_{p}$, it is known (see e.g., [KoShYa-2025, Lemma 2.2]) that
$\mbox{Ker}A_{p}=\{h\in C^{\infty}(\bar{\Omega});\,\operatorname{div}h=0,\,\,\operatorname{rot}h=0,\,\,\,h\cdot\nu|_{\partial\Omega}=0\}\quad\mbox{for all $1<p<\infty$}.$
Hence, in general $A_{2}$ is a non-negative self-adjoint operator in $L^{2}_{\sigma}(\Omega)$ with zero eigenvalue, and is strictly positive if and only if $\Omega$ is a simply connected domain in $\mathbb{R}^{3}$. However, by considering $\tilde{A}_{2}=A_{2}+1$ instead of $A_{2}$ itself, we may regard $A_{2}$ as a strictly positive self-adjoint operator in $L^{2}_{\sigma}(\Omega)$ with bounded inverse $A_{2}^{-1}$ in the proof of our main result. For such a correction, it suffices to replace $\varepsilon\Delta u_{\varepsilon}$ by $\varepsilon(-\Delta+1)u_{\varepsilon}$ in (1.2).
Let us define Besov spaces $B^{s}_{p,q}(A)$ in terms of $A=A_{2}$. Take a function $\phi\in C^{\infty}_{0}(-\infty,\infty)$ in such a way that
${\rm supp\,}\phi\in[2^{-1},2]\quad 0\leq\phi\leq 1,\quad\sum_{j\in\mathbb{Z}}\phi(2^{-j}\lambda)=1,\quad\lambda>0.$
We introduce $\{\phi_{j}\}_{j\in\mathbb{Z}}$ such that for every $j\in\mathbb{Z}$
$\phi_{j}(\lambda):=\phi(2^{-j}\lambda),\quad\lambda\in\mathbb{R},$ (1.3)
which is a partition of the unity by dyadic numbers on the half line $(0,\infty)$. For $1\leq p,q\leq\infty$ and $s\in\mathbb{R}$, the norm of $B^{s}_{p,q}(A)$ is given by
$\|u\|_{B^{s}_{p,q}(A)}:=\Big\|\Big(1-\sum_{j\geq 1}\phi_{j}(\sqrt{A})\Big)f\Big\|_{L^{p}}+\Big\|\Big\{2^{js}\|\phi_{j}(\sqrt{A})u\|_{L^{p}}\Big\}_{j\in\mathbb{N}}\Big\|_{\ell^{q}(\mathbb{N})}.$ (1.4)
Our main result reads as follows.
Theorem 1.1.
For every $u_{0}\in B^{\frac{5}{2}}_{2,1}(A)$ there exist a time $T>0$ and a unique solution $u\in C([0,T],B^{\frac{5}{2}}_{2,1}(A))$ of (1.1). Such a solution $u$ satisfies $u\in C^{1}([0,T],W^{1,p}(\Omega))$ for all $1\leq p\leq 3$.
Remark 1.2.
(i) The above theorem shows unique existence of (1.1) in the critical Besov space $B^{\frac{5}{2}}_{2,1}(A)$ in any bounded domain $\Omega\subset\mathbb{R}^{3}$ with smooth boundary $\partial\Omega$. It should be emphasized that any geometric restriction on $\Omega$ is unnecessary.
(ii) Obviously, we see that the corresponding result to Theorem 1.1 holds in $\mathbb{R}^{3}$. Bourgain and Li [BourLi-2021, Theorem 1.7] considered the Cauchy problem of the Euler equations in $\mathbb{R}^{n}$ for $n=2,3$, and proved ill-posedness in $B^{1+\frac{n}{p}}_{p,q}(\mathbb{R}^{n})$ for $1<p<\infty,1<q\leq\infty$. Hence, our result may be regarded as an optimal theorem in $\mathbb{R}^{3}$ which exhibits the borderline case between well-posedness and ill-posedness in Besov spaces.
(iii) Higher regularity yields a better approximation rate. Specifically, for $s=5/2$ and $u\in\dot{B}^{\frac{5}{2}}_{2,1}(A)$, there exists a constant $C>0$ such that
$\left\|u-\sum_{j\leq J}\phi_{j}(\sqrt{A})u\right\|_{\dot{B}^{\frac{5}{2}}_{2,1}(A)}\leq C2^{-\frac{5}{2}J}$
for all $J\in\mathbb{N}$. Furthermore, since the spectrum of the Stokes operator on a bounded domain consists only of eigenvalues, the partial sum above is a finite linear combination of the corresponding eigenfunctions.
This paper is organized as follows. In Section 2, we provide several preliminary lemmas and propositions required to prove Theorem 1.1. Section 3 is devoted to constructing solutions for the equation regularized by $\varepsilon\Delta$. In Section 4, we establish the uniform boundedness of the solutions obtained in Section 3 with respect to the parameter $\varepsilon>0$. Finally, in Section 5, we provide the proof of Theorem 1.1 by taking the limit as $\varepsilon\to 0$. Throughout the following sections, we suppose that $A_{2}$ is strictly positive.
2. Preliminary
In this section, we collect several preliminary definitions and results. In Subsection 2.1, we fundamental properties of the Stokes operator $A$ and the associated Besov spaces $B^{s}_{p,q}(A)$. Subsection 2.2 is devoted to providing several propositions and lemmas required for the proof of Theorem 1.1, including Sobolev-type embeddings, derivative estimates, and various properties of derivatives involving spectral restriction operators.
2.1. The Stokes operator and Besov spaces
We begin by recalling several fundamental results regarding vector fields, the Laplacian, and the Stokes operator (see [Miy-1980]). Following the procedure established in [IMT-2019, FI-2024], we introduce Besov and Sobolev spaces associated with the Stokes operator.
Let $1<p<\infty$. The Laplacian $B_{p}$ with the Neumann condition acting on vector fields in $L^{p}$ is defined as follows:
$\begin{cases}D(B_{p})=\{u\in W^{2,p}(\Omega)\,|\,{\rm rot}\,u\times\nu=0,u\cdot\nu=0\text{ on }\partial\Omega\},\\ B_{p}u=-\Delta u,\quad u\in D(B_{p}).\end{cases}$
For the Stokes operator $A_{p}$ on $L^{p}_{\sigma}$ defined in the previous section, we have $D(A_{p})=\mathbb{P}_{p}D(B_{p})$, and it holds that
$B_{p}\mathbb{P}_{p}=\mathbb{P}_{p}B=A_{p}\mathbb{P}_{p}\quad\mbox{on $D(B_{p})$}.$
We here explain the derivation of the projection $\mathbb{P}_{p}$ briefly. In the $L^{p}$ setting, the solvability of the following Neumann problem is well-established:
$\begin{cases}\Delta\pi=\mathrm{div}\,u&\text{in }\Omega,\\ \dfrac{\partial\pi}{\partial\nu}=u\cdot\nu&\text{on }\partial\Omega.\end{cases}$ (2.1)
More precisely, we consider the weak formulation of the Neumann problem. It is known that for a given vector field $u\in L^{p}$, there exists a unique gradient $\nabla\pi\in L^{p}$ such that
$\langle\nabla\pi,\nabla\phi\rangle=\langle u,\nabla\phi\rangle\quad\text{for all }\phi\in H^{1}_{p^{\prime}}(\Omega),$
where $\langle\cdot,\cdot\rangle$ denotes the dual coupling between $L^{p}$ and $L^{p^{\prime}}$ with $1/p+1/p^{\prime}=1$. The solvability of this problem is equivalent to the existence of a positive constant $C$ such that
$\|\nabla\pi\|_{L^{p}}\leq C\sup_{\phi\in H^{1}_{p^{\prime}}(\Omega)\setminus\{0\}}\frac{|\langle\nabla\pi,\nabla\phi\rangle|}{\|\nabla\phi\|_{L^{p^{\prime}}}}.$
Based on this, we define the Helmholtz projection $\mathbb{P}_{p}:L^{p}\to L^{p}_{\sigma}$ by
$\mathbb{P}_{p}u=u-\nabla\pi,$
where $\nabla\pi$ is the gradient of the solution to the weak Neumann problem described above. Furthermore, we note the following identity:
$\Delta\nabla\pi=\nabla\mathrm{div}\,u.$
For a smooth solution $(u,\pi)$ of the Euler equations (1.1), this identity yields
$\Delta\Big((u\cdot\nabla)u+\nabla\pi\Big)=\Delta\Big((u\cdot\nabla)u\Big)+\nabla\mathrm{div}\Big((u\cdot\nabla)u\Big).$ (2.2)
The following is a well-known fact regarding the orthogonality between divergence-free and rotation-free vector fields.
Lemma 2.1.
Let $1<p<\infty$ and $1/p+1/p^{\prime}=1$. For every scalar function $\pi\in W^{1,p^{\prime}}(\Omega)$ and every vector field $v\in L^{p}(\Omega)$, we have
$\int_{\Omega}\nabla\pi\cdot\mathbb{P}v\,dx=0.$
Proof.
For any $v\in L^{p}(\Omega)$, its Helmholtz projection satisfies $\mathbb{P}v\in L^{p}_{\sigma}(\Omega)$. By the density of $C^{\infty}_{0,\sigma}(\Omega)$ in $L^{p}_{\sigma}(\Omega)$ for $1<p<\infty$, there exists a sequence $\{u_{N}\}_{N=1}^{\infty}\subset C^{\infty}_{0,\sigma}(\Omega)$ such that $u_{N}\to\mathbb{P}v$ in $L^{p}(\Omega)$ as $N\to\infty$. For $\pi\in W^{1,p}(\Omega)$, it is straightforward that
$\int_{\Omega}\nabla p\cdot u_{N}~dx=0,$
and the equality for $\mathbb{P}v$ instead of $u_{n}$ is ensured by taking the limit as $N\to\infty$. ∎
The operator $A_{2}$ is a non-negative self-adjoint operator on $L^{2}_{\sigma}$. By the spectral theorem, there exists a resolution of the identity $\{E(\lambda)\}_{\lambda\in\mathbb{R}}$ such that
$u=\int_{-\infty}^{\infty}dE(\lambda)u\text{ in }L^{2}_{\sigma}\quad\text{for all }u\in L^{2}_{\sigma}.$
For any measurable function $\varphi:\mathbb{R}\to\mathbb{C}$, the operator $\varphi(A_{2})$ is well-defined on $L^{2}_{\sigma}$ as
$\varphi(A_{2})u=\int_{-\infty}^{\infty}\varphi(\lambda)dE(\lambda)u.$
The following proposition is fundamental for extending the framework of Besov and Sobolev spaces from $L^{2}$ to the $L^{p}$ setting.
Proposition $2.$
Let $1\leq p\leq\infty$, $\phi\in C_{0}^{\infty}(\mathbb{R})$, and $\phi_{j}(\lambda)=\phi(2^{-2j}\lambda)$ for $j\in\mathbb{Z}$ and $\lambda\in\mathbb{R}$. Then
$\sup_{j\in\mathbb{Z}}\|\phi_{j}(\sqrt{A_{2}})\|_{L^{p}\to L^{p}}<\infty.$
The proof of Proposition 2.2 is analogous to that in [IMT-2018] (see also Subsection 2.2 in [IwKo-preprint]). Next, we introduce the space of test functions and its dual space.
Definition. (Spaces of test functions and distributions)
  1. (1)
    We define the space of test functions $\mathcal{X}_{\sigma}$ by
    $\mathcal{X}_{\sigma}:=\{u\in L^{1}\cap L^{2}_{\sigma}\mid p_{M}(u)<\infty\text{ for all }M\in\mathbb{N}\},$
    where the semi-norms are given by
    $p_{M}(u):=\|u\|_{L^{1}}+\sup_{j\in\mathbb{N}}2^{Mj}\|\phi_{j}(A_{2})u\|_{L^{1}}.$
  2. (2)
    We denote by $\mathcal{X}_{\sigma}^{\prime}$ the topological dual space of $\mathcal{X}_{\sigma}$.
The spaces $\mathcal{X}_{\sigma}$ and $\mathcal{X}^{\prime}_{\sigma}$ are of non-homogeneous type. Since we may suppose that the operator $A_{2}$ is strictly positive (i.e., the infimum of its spectrum is strictly positive), the non-homogeneous and homogeneous spaces are equivalent. Here, we adopt the non-homogeneous framework.
We introduce the projection $\mathbb{P}$ onto the space of divergence-free vector fields and the Stokes operator on $\mathcal{X}_{\sigma}^{\prime}$. While the operator $A$ and its resolvent act on divergence-free fields, $B$ and its resolvent act on a more general class of vector fields.
Definition. (The projection $\mathbb{P}$, the Stokes operator $A$ and the Laplacian $B$ on $\mathcal{X}_{\sigma}^{\prime}$)
  1. (1)
    For $u\in L^{p}$ with $1\leq p\leq\infty$, we define $\mathbb{P}u\in\mathcal{X}_{\sigma}^{\prime}$ by
    $_{\mathcal{X}_{\sigma}^{\prime}}\langle\mathbb{P}u,v\rangle_{\mathcal{X}_{\sigma}}:=\int_{\Omega}u\cdot\mathbb{P}_{2}v\,dx\quad\text{for }v\in\mathcal{X}_{\sigma}.$
    Hereafter, we simply write $\mathbb{P}$ for $\mathbb{P}_{2}$.
  2. (2)
    For $u\in\mathcal{X}_{\sigma}^{\prime}$, we define $Au\in\mathcal{X}_{\sigma}^{\prime}$ by
    $_{\mathcal{X}_{\sigma}^{\prime}}\langle Au,v\rangle_{\mathcal{X}_{\sigma}}:={}_{\mathcal{X}_{\sigma}^{\prime}}\langle u,A_{2}v\rangle_{\mathcal{X}_{\sigma}}\quad\text{for }v\in\mathcal{X}_{\sigma}.$
  3. (3)
    Let $\phi\in C^{\infty}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$. For $u\in\mathcal{X}_{\sigma}^{\prime}$, we define $\phi(A)u\in\mathcal{X}_{\sigma}^{\prime}$ by
    $_{\mathcal{X}_{\sigma}^{\prime}}\langle\phi(A)u,v\rangle_{\mathcal{X}_{\sigma}}:={}_{\mathcal{X}_{\sigma}^{\prime}}\langle u,\phi(A_{2})v\rangle_{\mathcal{X}_{\sigma}}\quad\text{for }v\in\mathcal{X}_{\sigma}.$
    We write AA for $A_{2}$ when the context is clear.
  4. (4)
    Similarly, we define the operators $B$ and $\phi(B)$ on $\mathcal{X}_{\sigma}^{\prime}$ as the dual operators of $B_{2}$ and $\phi(B_{2})$, respectively. That is, for $u\in\mathcal{X}_{\sigma}^{\prime}$,
    $_{\mathcal{X}_{\sigma}^{\prime}}\langle Bu,v\rangle_{\mathcal{X}_{\sigma}}:={}_{\mathcal{X}_{\sigma}^{\prime}}\langle u,B_{2}v\rangle_{\mathcal{X}_{\sigma}},\quad{}_{\mathcal{X}_{\sigma}^{\prime}}\langle\phi(B)u,v\rangle_{\mathcal{X}_{\sigma}}:={}_{\mathcal{X}_{\sigma}^{\prime}}\langle u,\phi(B_{2})v\rangle_{\mathcal{X}_{\sigma}}$
    for all $v\in\mathcal{X}_{\sigma}$. We also abbreviate $B_{2}$ as $B$.
From the definition above, it follows that
$f=\mathbb{P}f\quad\text{in }\mathcal{X}^{\prime}_{\sigma},$
for any $f\in\mathcal{X}^{\prime}_{\sigma}\cap L^{p}$ with $1\leq p\leq\infty$.
The $L^{p}$-boundedness of the Helmholtz projection is well-established for $1<p<\infty$ (see, e.g., [Miy-1980, Section 2]). Since our operator $\mathbb{P}$ is defined as a natural extension to the distribution space $\mathcal{X}^{\prime}_{\sigma}$, its boundedness on $L^{p}$ follows immediately.
Lemma 2.3.
Let $1<p<\infty$. Then $\mathbb{P}$ is a bounded linear operator on $L^{p}$.
The following lemma shows that the resolvent of AA can be expressed using the resolvent of $B$ and the projection $\mathbb{P}$.
Lemma 2.4.
Let $1\leq p\leq\infty$ and $\mu\geq 0$. The operator $(\mu+A)^{-1}\mathbb{P}:L^{p}\to\mathcal{X}_{\sigma}^{\prime}$ is well-defined and satisfies
$(\mu+B)^{-1}\mathbb{P}=\mathbb{P}(\mu+B)^{-1}=(\mu+A)^{-1}\mathbb{P}.$ (2.3)
Remark. This lemma is intended to be applied to the nonlinear term $(u\cdot\nabla)u$, which in general does not satisfy the divergence-free condition or the prescribed boundary conditions. The identity (2.3) also holds as an operator identity on $\mathcal{X}^{\prime}_{\sigma}$.
Proof. It is shown in [Miy-1980, Corollary 3.6] that the identity
$(\mu+B_{p})^{-1}\mathbb{P}_{p}=\mathbb{P}_{p}(\mu+B_{p})^{-1}=(\mu+A_{p})^{-1}\mathbb{P}_{p}$
holds for any $\mu$ in the resolvent set of $B_{p}$ within the $L^{p}$ framework for $1<p<\infty$. By the definition of the projection $\mathbb{P}:L^{p}\to\mathcal{X}^{\prime}_{\sigma}$ and the duality argument, we obtain the well-definedness and the desired identity (2.3). $\Box$
We now define the Besov and Sobolev spaces associated with the Stokes operator AA.
Definition. (Besov and Sobolev spaces)  Let $s\in\mathbb{R}$ and $1\leq p,q\leq\infty$. We define the Besov space $B^{s}_{p,q}(A)$ associated with the Stokes operator AA by
$B^{s}_{p,q}(A):=\left\{u\in\mathcal{X}_{\sigma}^{\prime}\Bigm|\|u\|_{B^{s}_{p,q}(A)}:=\left\|\left\{2^{sj}\|\phi_{j}(\sqrt{A})u\|_{L^{p}}\right\}_{j\in\mathbb{Z}}\right\|_{\ell^{q}(\mathbb{Z})}<\infty\right\},$
where $\{\phi_{j}\}_{j\in\mathbb{Z}}$ is a dyadic partition of unity satisfying (1.3). Similarly, for $s\in\mathbb{R}$, we define the Sobolev space $H^{s}(A)$ by
$H^{s}(A):=\left\{u\in\mathcal{X}_{\sigma}^{\prime}\Bigm|\|u\|_{H^{s}(A)}:=\|A^{\frac{s}{2}}u\|_{L^{2}}<\infty\right\}.$
Remark. In the definition above, we adopt the norm of homogeneous type as an equivalent norm for $B^{s}_{p,q}(A)$. We note that the homogeneous and non-homogeneous Besov spaces are equivalent under the present assumptions on $\Omega$, since the infimum of the spectrum of the Stokes operator AA is strictly positive. Consequently, the non-homogeneous norm introduced in (1.4) is equivalent to the homogeneous one as follows:
$\begin{split}&\left\|\left\{2^{js}\|\phi_{j}(\sqrt{A})u\|_{L^{p}}\right\}_{j\in\mathbb{Z}}\right\|_{\ell^{q}(\mathbb{Z})}.\\ &\simeq\left\|\left(1-\sum_{j=1}^{\infty}\phi_{j}(\sqrt{A})\right)u\right\|_{L^{p}}+\left\|\left\{2^{js}\|\phi_{j}(\sqrt{A})u\|_{L^{p}}\right\}_{j=1}^{\infty}\right\|_{\ell^{q}(\mathbb{N})}.\end{split}$
Definition. We denote $D:=\sqrt{A},\sqrt{B}$.
Remark. We frequently employ the notation $D$ for spectral restriction operators instead of $\sqrt{A}$ or $\sqrt{B}$. This is justified by the fact that
$\phi_{j}(\sqrt{A})u=\phi_{j}(\sqrt{B})\mathbb{P}u=\phi_{j}(\sqrt{B})u\quad\text{in }L^{p}_{\sigma}$
for $u\in L^{p}_{\sigma}$. The divergence-free property is crucial for this identity. Indeed, if the divergence-free condition is not assumed, $\phi_{j}(\sqrt{B})\mathbb{P}u$ remains well-defined for any $u\in L^{p}$, whereas $\phi_{j}(\sqrt{A})u$ is generally not defined for arbitrary vector fields. There is an essential difference in the boundary conditions inherent in the definitions of AA and $B$; specifically, operators associated with AA act on vector fields that intrinsically satisfy the divergence-free condition. In our subsequent analysis, the nonlinear term $(u\cdot\nabla)u$ will be handled by applying the projection $\mathbb{P}$.
$2.$ Analytical tools
In this subsection, we provide several analytical tools required for our main results. These include elliptic estimates involving weak derivatives and the operators AA and $B$, as well as fundamental properties of spectral restriction operators and the Stokes semigroup.
Lemma 2.5.
Let $1<p<\infty$.
  1. (1)
    (Elliptic estimates) There exists a constant $C>0$ such that for all $u\in L^{p}_{\sigma}$,
    $\|\nabla^{2}A^{-1}u\|_{L^{p}}\leq C\|u\|_{L^{p}}.$
  2. (2)
    There exists a constant $C>0$ such that for all $u\in L^{p}_{\sigma}$,
    $\|\nabla A^{-\frac{1}{2}}u\|_{L^{p}}\leq C\|u\|_{L^{p}}.$
  3. (3)
    The operators $\nabla^{2}B^{-1}$ and $\nabla B^{-\frac{1}{2}}$ are bounded linear operators on $L^{p}$.
  4. (4)
    For any $u\in L^{2}$, the identity $A^{\frac{1}{2}}\mathbb{P}u=A\mathbb{P}B^{-\frac{1}{2}}u$ holds in $\mathcal{X}^{\prime}_{\sigma}$.
Proof.
(1) and (2) are established in [IwKo-preprint, Proposition 2.1]. (3) can be proved by an argument analogous to (1) and (2). For (4), we recall the identity $A_{2}^{\frac{1}{2}}u=A_{2}A_{2}^{-\frac{1}{2}}u$ in $L^{2}$ for all $u\in H^{\frac{1}{2}}(A)$. We utilize the integral representation of the fractional power:
$A_{2}^{-\frac{1}{2}}=c_{0}\int_{0}^{\infty}t^{-\frac{1}{2}}(1+tA_{2})^{-1}dt,$
where $c_{0}$ is an appropriate positive constant. This representation, together with the commutativity relation (2.3), yields that for $u\in L^{2}$
$A_{2}A_{2}^{-\frac{1}{2}}u=A_{2}c_{0}\int_{0}^{\infty}t^{-\frac{1}{2}}(1+tA_{2})^{-1}u~dt=A_{2}\mathbb{P}c_{0}\int_{0}^{\infty}t^{-\frac{1}{2}}(1+tB_{2})^{-1}u~dt=A_{2}\mathbb{P}B_{2}^{-\frac{1}{2}}u.$
The identity (4) then follows from the definitions of AA and $B$ on $\mathcal{X}^{\prime}_{\sigma}$. ∎
Lemma 2.6.
Let $m\in\mathbb{N}$ and $1<p<\infty$. Then a positive constant $C$ exists such that
$\|B^{-1}u\|_{W^{m+2,p}(\Omega)}\leq C\|u\|_{W^{m,p}(\Omega)}\quad\text{for }u\in W^{m,p}(\Omega).$
Proof. The case when $m=1,2$ is already established in Lemma 2.5. When $m\geq 3$, we apply Lemma 4.4 in [KoYa-2009] under the boundary condition $u\cdot\nu={\rm rot\,}u\times\nu=0$. $\Box$
Lemma 2.7.
Let $m\in\mathbb{N}$ and $1<p<\infty$. Then a positive constant $C$ exists such that
[…]
Proof. Following the proof of Lemma 3.3 (i) in [GiMi-1985], the above inequality is proved. In fact, we write $\mathbb{P}u=u-\nabla\pi$, where $\pi$ is the solution of the Neumann problem (2.1), and the elliptic esimate in $W^{m,p}(\Omega)$ for $\nabla\pi$ is available. $\Box$
Lemma 2.8.
  1. (1)
    We have the norm equivalence that for every $u\in H^{1}(A)$
    $\|u\|_{H^{1}(A)}\simeq\|\nabla u\|_{L^{2}}.$
    Furthermore, for every $u\in L^{2}$ with $\|B^{\frac{1}{2}}u\|_{L^{2}}<\infty$,
    $\|B^{\frac{1}{2}}u\|_{L^{2}}\simeq\|\nabla u\|_{L^{2}}.$
  2. (2)
    Let |s2|<1|s-2|<1. We have the norm equivalece that for every $u\in B^{s}_{2,1}(A)$
    $\|u\|_{B^{s}_{2,1}(A)}\simeq\sum_{j\in\mathbb{Z}}2^{(s-2)j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u\|_{L^{2}},$
    where $\psi_{j}$ is defined by
$\psi_{j}(\lambda)=(1+2^{-2j-2}\lambda^{2})^{-1}-(1+2^{-2j}\lambda^{2})^{-1},\quad\lambda\geq 0.$ (2.4)
Remark. The norm on the right-hand side of Lemma 2.8 (2) does not directly control the boundary values of $u$; however, it imposes a restriction on $\Delta u$.
Proof. (1) Let $u\in H^{1}(A)$ and define $u_{N}=\displaystyle\sum_{j\leq N}\phi_{j}(D)u$. It is straightforward to see that $u_{N}$ converges to $u$ in $H^{1}(A)$ as $N\to\infty$, and that $u_{N}\in H^{2}(A)$. By using $\Delta=-{\rm rot\,rot\,}+\nabla{\rm div}$, and the fact that ${\rm div\,}u_{N}=0\text{ in }\Omega$ and ${\rm rot\,}u_{N}\times\nu=0\text{ on }\partial\Omega$ , we have
$\|u_{N}\|_{H^{1}(A)}=\int_{\Omega}(-\Delta u_{N})\cdot u_{N}~dx=\int_{\Omega}({\rm rot\,rot\,}u_{N})\cdot u_{N}~dx=\|{\rm rot\,}u_{N}\|_{L^{2}}^{2}\leq C\|\nabla u_{N}\|_{L^{2}}^{2},$
which proves one side of the inequality. Conversely, we apply in [KoYa-2009, Theorem 2.4] to obtain the reverse inequality. Under the assumption that $\Omega$ excludes non-trivial harmonic fields, we have
$\|\nabla u_{N}\|_{L^{2}}\leq C\|{\rm rot\,}u_{N}\|_{L^{2}}+C\|{\rm div\,}u_{N}\|_{L^{2}}=C\|u_{N}\|_{H^{1}(A)},$
where we have used the divergence-free property of $u_{N}$. Thus $\{u_{N}\}_{N\in\mathbb{N}}$ is a bounded sequence in $H^{1}(\Omega)$ and converges to $u$ strongly in $L^{2}(\Omega)$. It follows that the limit $u$ belongs to $H^{1}(\Omega)\cap L^{2}_{\sigma}$. Taking the limit in the equivalence $\|u_{N}\|_{H^{1}(A)}\simeq\|\nabla u_{N}\|_{L^{2}}$ as $N\to\infty$, we conclude that
$\|u\|_{H^{1}(A)}\simeq\|\nabla u\|_{L^{2}}.$
(2) First, we observe that $\phi_{j}(\lambda)\psi_{j}(\lambda)^{-\frac{1}{2}}$ for $\lambda\in\mathbb{R}$ is a smooth function with compact support. Furthermore, it satisfies the scaling property
$\phi_{j}(\lambda)\psi_{j}(\lambda)^{-\frac{1}{2}}=\phi_{0}(2^{-j}\lambda)\psi_{0}(2^{-j}\lambda)^{-\frac{1}{2}}.$
By the spectral theorem, we have the uniform boundedness of $\phi_{j}(D)\psi_{j}(D)^{-\frac{1}{2}}$ with respect to $j\in\mathbb{Z}$ as operators on $L^{2}$. We observe that the operator $\phi_{j}(D)$ restricts the spectrum to a dyadic shell around $2^{j}$, where the operator $B$ is bounded from below by $C2^{2j}$. Consequently, we have
$2^{2j}\|\phi_{j}(D)u\|_{L^{2}}\leq C\|B\phi_{j}(D)\psi_{j}(D)^{-\frac{1}{2}}\psi_{j}(D)u\|_{L^{2}}\leq C\|\psi_{j}(D)^{\frac{1}{2}}\Delta u\|_{L^{2}}.$
Multiplying by $2^{(s-2)j}$ and summing over $j\in\mathbb{Z}$, we arrive at
$\|u\|_{\dot{B}^{s}_{2,1}(A)}\leq C\sum_{j\in\mathbb{Z}}2^{(s-2)j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u\|_{L^{2}}.$
Conversely, we prove the reverse inequality with the decomposition by $\{\phi_{j}\}_{j\in\mathbb{Z}}$. We note that
$\psi_{j}(D)^{\frac{1}{2}}=\bigg(\frac{3}{4}\cdot 2^{-2j}B(1+2^{(-2j-2)}B)^{-1}(1+2^{-2j}B)^{-1}\bigg)^{\frac{1}{2}}\sum_{k\in\mathbb{Z}}\phi_{k}(D)$
and for each $k$, it follows that
$\begin{split}&\bigg\|\bigg(\frac{3}{4}\cdot 2^{-2j}B(1+2^{(-2j-2)}B)^{-1}(1+2^{-2j}B)^{-1}\bigg)^{\frac{1}{2}}\phi_{k}(D)\Delta u\bigg\|_{L^{2}}\\ \leq&C\dfrac{2^{-j+k}}{1+2^{-2j+2k}}\|\phi_{k}(D)\Delta u\|_{L^{2}}\\ \leq&C2^{-|j-k|}\|\phi_{k}(D)\Delta u\|_{L^{2}}.\end{split}$
Thus, we have
$\begin{split}\sum_{j\in\mathbb{Z}}2^{(s-2)j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u\|_{L^{2}}\leq&C\sum_{j\in\mathbb{Z}}\sum_{k\in\mathbb{Z}}2^{(s-2)(j-k)}2^{-|j-k|}\cdot 2^{(s-2)k}\|\phi_{k}(D)\Delta u\|_{L^{2}}\\ \leq&C\Big(\sum_{j\in\mathbb{Z}}2^{(s-2)j-|j|}\Big)\sum_{k\in\mathbb{Z}}2^{(s-2)k}\|\phi_{k}(D)\Delta u\|_{L^{2}}\\ \leq&C\|u\|_{B^{s}_{2,1}(A)}.\end{split}$
Since |s2|<1|s-2|<1 ensures the convergence of the series $\sum_{m\in\mathbb{Z}}2^{(s-2)m-|m|}$, the proof is complete. $\Box$
Lemma 2.9.
Let $\psi_{j}$ be defined by (2.4). Then for every $j\in\mathbb{Z}$
$\|\psi_{j}(D)\|_{L^{2}\to L^{2}}\leq\|\psi_{j}(D)^{\frac{1}{2}}\|_{L^{2}\to L^{2}}.$ (2.5)
A positive constant $C$ exists such that for every $j\in\mathbb{Z}$
$\|\nabla(1+2^{-2j}B)^{-1}\|_{L^{2}\to L^{2}}\leq C2^{j}\|\psi_{j}(D)^{\frac{1}{2}}\|_{L^{2}\to L^{2}}.$ (2.6)
Proof. By the definition of $\psi_{j}(D)$,
$\psi_{j}(D)=\dfrac{3}{4}\cdot 2^{-2j}B(1+2^{-2j-2}B)^{-1}(1+2^{-2j}B)^{-1},$
is an non-negative operator on $L^{2}$, and we can write $\psi_{j}(D)=\big(\psi_{j}(D)^{\frac{1}{2}}\big)^{2}$ with the estimate $\|\psi_{j}(D)^{\frac{1}{2}}\|_{L^{2}\to L^{2}}\leq 1$, which proves (2.5).
By Lemma 2.8 (1),
$\|\nabla(1+2^{-2j}B)^{-1}\|_{L^{2}\to L^{2}}\leq C\|B^{\frac{1}{2}}(1+2^{-2j}B)^{-1}\|_{L^{2}\to L^{2}},$
and we write
$B^{\frac{1}{2}}(1+2^{-2j}B)^{-1}=\bigg[\Big(\frac{3}{4}(1+2^{(-2j-2)}B)^{-1}(1+2^{-2j}B)^{-1}\Big)^{-\frac{1}{2}}(1+2^{-2j}B)^{-1}\bigg]2^{j}\psi_{j}(D)^{\frac{1}{2}}.$
Since the right hand side without $2^{j}\psi_{j}(D)^{\frac{1}{2}}$ is bounded uniformly with respect to $j\in\mathbb{Z}$, we obtain the inequality (2.6). $\Box$
Lemma 2.10.
(Proposition 2.3 in [IwKo-preprint]) A positive constant $C$ exists such that for every $1\leq p\leq r\leq\infty$ and $j\in\mathbb{Z}$
$\|\phi_{j}(D)\|_{L^{r}\to L^{p}}\leq C2^{3(\frac{1}{r}-\frac{1}{p})j}.$ (2.7)
Lemma 2.11.
(Embedding in Besov spaces) Let $s>0$, $1\leq r\leq p\leq\infty$ and $1\leq q\leq\infty$. Then
$\begin{split}\|u\|_{B^{0}_{p,q}(A)}\leq&C\|u\|_{B^{3(\frac{1}{r}-\frac{1}{p})}_{r,q}(A)},\quad u\in B^{3(\frac{1}{r}-\frac{1}{p})}_{r,q}(A),\\ \|u\|_{B^{0}_{p,1}(A)}\leq&C\|u\|_{B^{s}_{p,\infty}(A)},\quad u\in B^{s}_{p,\infty}(A).\end{split}$
Furthermore, if $1\leq r\leq p<\infty$ or $1\leq r<p=\infty$, then
$\|\nabla u\|_{L^{p}}\leq C\|u\|_{B^{1+3(\frac{1}{r}-\frac{1}{p})}_{r,1}(A)},\quad u\in B^{1+3(\frac{1}{r}-\frac{1}{p})}_{r,1}(A).$
Proof. The proofs of the first and second inequalities are analogous to those of [IMT-2019, Proposition 3.2], based on the $L^{r}$-$L^{p}$ boundedness in (2.7) and the positivity of the spectrum.
For the third inequality, we utilize the resolution of the identity:
$u=\sum_{j\in\mathbb{Z}}\phi_{j}(D)u,$
which is justified by an argument similar to that in [IMT-2019, Lemma 4.5] for the Dirichlet Laplacian. Let $\Phi_{j}(D):=\phi_{j-1}(D)+\phi_{j}(D)+\phi_{j+1}(D)$. Then we have the property $\phi_{j}(D)=\Phi_{j}(D)\phi_{j}(D)$. Moreover, if $p<\infty$, then the following $L^{r}$-$L^{p}$ boundedness holds:
$\|\nabla\Phi_{j}(D)\|_{L^{r}\to L^{p}}\leq C\|B^{\frac{1}{2}}\Phi_{j}(D)\|_{L^{r}\to L^{p}}\leq C2^{j+3\left(\frac{1}{r}-\frac{1}{p}\right)j}.$
Applying these estimates, we obtain
$\|\nabla u\|_{L^{p}}\leq\sum_{j\in\mathbb{Z}}\|\nabla\Phi_{j}(D)\phi_{j}(D)u\|_{L^{p}}\leq C\sum_{j\in\mathbb{Z}}2^{j+3\left(\frac{1}{r}-\frac{1}{p}\right)j}\|\phi_{j}(D)u\|_{L^{p}},$
which proves the proof of the third inequality for $p<\infty$. When $p=\infty$, we utilize the interpolation inequality (see [Niren-1959]) and the elliptic estimate. Let $\max\{r,n\}<r_{0}<\infty$. Then
$\begin{split}\|\nabla\Phi_{j}(D)\phi_{j}(D)u\|_{L^{\infty}}&\leq C\|\nabla\Phi_{j}(D)\phi_{j}(D)u\|_{L^{r_{0}}}^{1-\frac{3}{r_{0}}}\|\nabla^{2}\Phi_{j}(D)\phi_{j}(D)u\|_{L^{r_{0}}}^{\frac{3}{r_{0}}}\\ &\leq C\|B^{\frac{1}{2}}\Phi_{j}(D)\phi_{j}(D)u\|_{L^{r_{0}}}^{1-\frac{3}{r_{0}}}\|B\Phi_{j}(D)\phi_{j}(D)u\|_{L^{r_{0}}}^{\frac{3}{r_{0}}}\\ &\leq C2^{(1+\frac{3}{r_{0}})j}\|\phi_{j}(D)u\|_{L^{r_{0}}}\\ &\leq C2^{(1+\frac{3}{r})j}\|\phi_{j}(D)u\|_{L^{r}},\end{split}$
which completes the proof for $p=\infty$. $\Box$
Lemma 2.12.
(Sobolev embedding) Let $1<r\leq p<\infty$. Then
$\|u\|_{L^{p}}\leq C\|A^{\frac{3}{2}(\frac{1}{r}-\frac{1}{p})}u\|_{L^{r}},\quad\text{for }u\text{ with }\|A^{\frac{3}{2}(\frac{1}{r}-\frac{1}{p})}u\|_{L^{r}}<\infty.$
Let $1<p<\infty$ and $s\geq 1+3(1/r-1/p)$. Then
$\|\nabla u\|_{L^{p}}\leq C\|A^{\frac{s}{2}}u\|_{L^{r}},\quad\text{for }u\text{ with }\|A^{\frac{s}{2}}u\|_{L^{r}}<\infty.$
Proof. The first inequality can be proved in a similar manner to [FI-2024, Theorem 1.4]. For the second inequality, we apply Lemma 2.5 and the first inequality to obtain
$\|\nabla u\|_{L^{p}}\leq C\|A^{\frac{1}{2}}u\|_{L^{p}}\leq C\|A^{\frac{1}{2}+\frac{3}{2}\left(\frac{1}{r}-\frac{1}{p}\right)}u\|_{L^{r}}.$
An embedding between regularity indices, ensured by the positivity of the spectrum, yields
$\|A^{\frac{1}{2}+\frac{3}{2}\left(\frac{1}{r}-\frac{1}{p}\right)}u\|_{L^{r}}\leq C\|A^{\frac{s}{2}}u\|_{L^{r}},$
which completes the proof. $\Box$
Proposition 2.13.
Let $s>0$ and $1\leq r\leq p\leq\infty$. Then for $0<t<1$
$\begin{split}\|e^{-tA}u_{0}\|_{L^{p}}\leq&Ct^{-\frac{3}{2}(\frac{1}{r}-\frac{1}{p})}\|u_{0}\|_{L^{r}},\quad\text{ for }u\in\mathcal{X}^{\prime}_{\sigma}\cap L^{r},\\ \|A^{s}e^{-tA}u_{0}\|_{L^{p}}\leq&Ct^{-\frac{s}{2}}\|u_{0}\|_{L^{p}},\quad\text{for }u\in\mathcal{X}^{\prime}_{\sigma}\cap L^{r}.\end{split}$
Remark. Proposition 2.13 provides local-in-time estimates. While global-in-time estimates with exponential decay could be expected, the present paper focuses primarily on the local-in-time analysis.
Proof. We know that the semigroup $e^{-tA}$ satsifies the Gaussian upper bound (see Lemma 2.2 [IwKo-preprint]), and the proof of Proposition 2.13 is similar to Theorem 1.6 (ii) in [FI-2024]. $\Box$
Proposition 2.14.
(Maximal regularity) Let $T>0$. A positive constant $C$, independent of $T$, exists such that for every $u_{0}\in H^{1}(A)$
$\|e^{-tA}u_{0}\|_{L^{2}(0,T;H^{2}(A))}\leq C\|u_{0}\|_{H^{1}(A)}.$
For every $f\in L^{2}(0,T;H^{2}(A))$
$\Big\|\int_{0}^{t}e^{-(t-\tau)A}f(\tau)~d\tau\Big\|_{L^{2}(0,T;H^{2}(A))}\leq C\|f\|_{L^{2}(0,T;H^{2}(A))}.$
Proof. The proof is similar to those of Theorems 1.3 and 1.4 in [Iw-2018]. $\Box$
Lemma 2.15.
Let $j\in\mathbb{Z}$. For $u\in L^{2}_{\sigma}$ with $\Delta u\in L^{2}$, we consider the linear operator defined by $u\mapsto\psi_{j}(D)\Delta u$. Then $\psi_{j}(D)\Delta$ is extended uniquely to a bounded linear operator on $L^{2}_{\sigma}$ and
$\sup_{j\in\mathbb{Z}}2^{-2j}\|\psi_{j}(D)\Delta\|_{L^{2}\to L^{2}}<\infty,\quad\int_{\Omega}\Big(\psi_{j}(D)(-\Delta u)\Big)\cdot u~dx\geq 0.$
Proof. For every $u\in C_{0,\sigma}^{\infty}$, the identity $\psi_{j}(D)(-\Delta)u=\psi_{j}(D)Au$ holds. We may write
$\psi_{j}(D)A=\frac{3}{4}(2^{-2j}A)^{2}(1+2^{-2j-2}A)^{-1}(1+2^{-2j}A)^{-1}\cdot 2^{2j},$
which implies the estimate
$\|\psi_{j}(D)Au\|_{L^{2}}\leq C2^{2j}\|u\|_{L^{2}}.$
This boundedness allows us to extend $\psi_{j}(D)A$ uniquely to a bounded linear operator on $L^{2}_{\sigma}$, since $C_{0,\sigma}^{\infty}$ is dense in $L^{2}_{\sigma}$.
Now, suppose $u\in L^{2}_{\sigma}$ and let $\{u_{n}\}\subset C_{0,\sigma}^{\infty}$ be a sequence such that $u_{n}\to u$ in $L^{2}$ as $n\to\infty$. For each $n$, it follows from the self-adjointness of the operator $\psi_{j}(D)$ and the properties of AA that
$\int_{\Omega}\left(\psi_{j}(D)Au_{n}\right)\cdot u_{n}\,dx=\int_{\Omega}|\psi_{j}(D)^{\frac{1}{2}}A^{\frac{1}{2}}u_{n}|^{2}\,dx\geq 0.$
By taking the limit as $n\to\infty$, we obtain
$\int_{\Omega}\left(\psi_{j}(D)Au\right)\cdot u\,dx\geq 0,$
which completes the proof. $\Box$
3. Solutions with $\varepsilon\Delta$ and Regularity
The goal of this section is to construct solutions to the Navier-Stokes equations with viscosity $\varepsilon\Delta$ for each parameter $\varepsilon>0$ in the space $C([0,T];H^{1}(A))$, supplemented with additional regularity for $\partial_{t}u$ and $\nabla^{4}u$. To this end, we provide two propositions: Proposition 3.1 ensures the existence of a local-in-time solution, while Proposition 3.2 establishes the higher regularity required to justify the $L^{2}$ energy estimates with localized spectra, which are essential for the proof of the main theorem.
Proposition 3.1.
Let $\varepsilon>0$ be given. Suppose that the initial data $u_{0}$ belongs to the $H^{1}(A)$ Sobolev space associated with the Stokes operator. Then there exist a time $T>0$ and a unique solution $u\in C([0,T],H^{1}(A))$ to the following integral equation:
$u(t)=e^{-t\varepsilon A}u_{0}-\int_{0}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}\Big((u(\tau)\cdot\nabla)u(\tau)\Big)d\tau,\quad t\in[0,T].$ (3.1)
Proof. We begin by establishing the existence of a local solution $u$. We first choose a constant $M_{0}$ satisfying
$\|u_{0}\|_{H^{1}(A)}\leq M_{0}.$
Let $T>0$ be a time to be specified later. We apply the Banach fixed point theorem to the complete metric space $X_{T}$, which is defined as
$X_{T}:=\left\{u\in C([0,T],H^{1}(A))\,\middle|\,\|u\|_{X_{T}}\leq 2M_{0}\right\},$
equipped with the norm
$\|u\|_{X_{T}}:=\sup_{t\in[0,T]}\|u(t)\|_{H^{1}(A)},$
and the metric
$d(u,v):=\|u-v\|_{X_{T}}.$
It is straightforward to see that
$\|e^{-t\varepsilon A}u_{0}\|_{X_{T}}\leq\|u_{0}\|_{H^{1}(A)}\leq M_{0}.$
We apply the smoothing property and the $L^{3/2}-L^{2}$ estimate in Proposition 2.13 for the semigroup $e^{-(t-\tau)\varepsilon A}$, and subsequently use the boundedness of the projection $\mathbb{P}$ on $L^{3/2}$ from Lemma 2.3 and the Hölder inequality. The following estimate holds:
$\begin{split}\Big\|\int_{0}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}\big((u\cdot\nabla)u\big)d\tau\Big\|_{H^{1}(A)}\leq&C\int_{0}^{t}\Big(\varepsilon(t-\tau)\Big)^{-\frac{1}{2}-\frac{3}{2}(\frac{2}{3}-\frac{1}{2})}\|\mathbb{P}\big((u\cdot\nabla)u\big)\|_{L^{\frac{3}{2}}}d\tau\\ \leq&C\varepsilon^{-\frac{3}{4}}\int_{0}^{t}(t-\tau)^{-\frac{3}{4}}\|u\|_{L^{6}}\|\nabla u\|_{L^{2}}d\tau.\end{split}$ (3.2)
The Sobolev embedding $H^{1}(A)\hookrightarrow L^{6}$ from Lemma 2.12 and the norm equivalence $\|\nabla u\|_{L^{2}}\simeq\|u\|_{H^{1}(A)}$ in Lemma 2.8 yield the following bound:
$\begin{split}\Big\|\int_{0}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}\big((u\cdot\nabla)u\big)d\tau\Big\|_{H^{1}(A)}\leq&C\varepsilon^{-\frac{3}{4}}T^{\frac{1}{4}}\|u\|_{X_{T}}^{2}\leq C\varepsilon^{-\frac{3}{4}}T^{\frac{1}{4}}M_{0}^{2},\end{split}$ (3.3)
for any $u\in X_{T}$. Analogously, we obtain the difference estimate:
$\begin{split}&\Big\|\int_{0}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}\big((u\cdot\nabla)u\big)d\tau-\int_{0}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}\big((v\cdot\nabla)v\big)d\tau\Big\|_{\dot{H}^{1}(A)}\\ \leq&C\varepsilon^{-\frac{3}{4}}T^{\frac{1}{4}}(\|u\|_{X_{T}}+\|v\|_{X_{T}})\|u-v\|_{X_{T}}\\ \leq&C\varepsilon^{-\frac{3}{4}}T^{\frac{1}{4}}M_{0}\|u-v\|_{X_{T}},\end{split}$ (3.4)
for any $u,v\in X_{T}$. By these inequalities and by choosing $T$ such that $C\varepsilon^{-\frac{3}{4}}T^{\frac{1}{4}}\ll 1$, we ensure that the map
$X_{T}\ni u\quad\mapsto\quad e^{-t\varepsilon A}u_{0}-\int_{0}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}\Big((u\cdot\nabla)u\Big)~d\tau\in X_{T}$
is well-defined and contractive. We then find a unique fixed point $u$ in $X_{T}$, which is the desired solution.
We now prove the uniqueness of the solution in the space $C([0,T],H^{1}(A))$. Suppose that $u$ and $v$ are two solutions of the integral equation with the same initial data $u_{0}$ in $H^{1}(A)$. We define the quantity
$M:=\sup_{t\in[0,T]}(\|u\|_{X_{T}}+\|v\|_{X_{T}}).$
Let $0<t_{0}<T$ and consider the difference uvu-v on the time interval $0,t_{0}]$. A similar argument to that used to derive inequality (3.4) implies that
$\|u-v\|_{X_{t_{0}}}\leq C\varepsilon^{-\frac{3}{4}}t_{0}^{\frac{1}{4}}M\|u-v\|_{X_{t_{0}}}.$
By choosing $t_{0}$ sufficiently small such that $C\varepsilon^{-\frac{3}{4}}t_{0}^{\frac{1}{4}}<1$, the above inequality yields $\|u-v\|_{X_{t_{0}}}=0$, which means $u=v$ on the time interval $0,t_{0}]$. Due to the time continuity, we have $u(t_{0})=v(t_{0})$ in $H^{1}(A)$. Treating $u(t_{0})$ as the new initial data, we can apply the same argument iteratively on the time interval $t_{0},2t_{0}]$, provided $2t_{0}\leq T$. By repeating this procedure finitely many times, the uniqueness is established throughout the entire interval $[0,T]$. $\Box$
We will need the regularity of the solutions to justify the energy estimate in the proof of our theorem. An analogous argument to Giga and Miyakawa [GiMi-1985] combined with maximal regularity ensures the following regularity.
Proposition 3.2.
Let $u\in C([0,T],H^{1}(A))$ be the unique local solution to the integral equation (3.1) constructed in the previous proposition. Then,
$u\in L^{2}(0,T;H^{3}(\Omega))\cap L^{2}(0,T;H^{2}(A)),$
and for any $t_{0}$ such that $0<t_{0}<T$, the solution $u$ satisfies the following higher regularity estimates:
$u\in L^{2}(t_{0},T;H^{4}(\Omega))\quad\text{and}\quad\partial_{t}u\in L^{2}(t_{0},T;H^{2}(A)).$
Here, $H^{2}(A)$ denotes the Sobolev space associated with the Neumann Stokes operator $A$, and $H^{4}(\Omega)$ is the standard Sobolev space equipped with the domain $\Omega$ (without boundary conditions).
Proof. Step 1. We first prove that the solution $u$ belongs to $L^{\infty}_{\rm loc}((0,T],H^{s}(A))$ for any $s<2$. If $1<s_{1}<3/2$, a similar argument to those used in deriving (3.2) and (3.3) shows that
$\begin{split}\|u(t)\|_{H^{s_{1}}(A)}\leq&Ct^{-\frac{s_{1}-1}{2}}\|u_{0}\|_{H^{1}(A)}+C\int_{0}^{t}\Big(\varepsilon(t-\tau)\Big)^{-\frac{s_{1}}{2}-\frac{3}{2}(\frac{2}{3}-\frac{1}{2})}\|\mathbb{P}(u\cdot\nabla)u\|_{L^{\frac{3}{2}}}~d\tau\\ \leq&Ct^{-\frac{s_{1}-1}{2}}\|u_{0}\|_{H^{1}(A)}+C\varepsilon^{-\frac{s_{1}}{2}-\frac{1}{4}}t^{-\frac{s_{1}}{2}+\frac{3}{4}}\sup_{\tau\in[0,t]}\|u(\tau)\|_{H^{1}(A)}^{2},\end{split}$
which implies that $u\in L^{\infty}_{\rm loc}((0,T],H^{s_{1}}(A))$.
Next, we choose indices $s_{2},p_{1},p_{2}$ satisfying the conditions
$\frac{3}{2}\leq s_{2}<2,,\quad 6<p_{1}<\infty,\quad 2<p_{2}<3,\quad\frac{1}{2}=\frac{1}{p_{1}}+\frac{1}{p_{2}},$
and it holds that
$\begin{split}\|u(t)\|_{H^{s_{2}}(A)}\leq&Ct^{-\frac{s_{2}-1}{2}}\|u_{0}\|_{H^{1}(A)}+C\int_{0}^{t}\Big(\varepsilon(t-\tau)\Big)^{-\frac{s_{2}}{2}}\|\mathbb{P}(u\cdot\nabla)u\|_{L^{2}}~d\tau\\ \leq&Ct^{-\frac{s_{2}-1}{2}}\|u_{0}\|_{H^{1}(A)}+C\varepsilon^{-\frac{s_{2}}{2}}t^{1-\frac{s_{2}}{2}}\sup_{\tau\in[0,t]}\|u\|_{L^{p_{1}}}\|\nabla u\|_{L^{p_{2}}}.\end{split}$
Now, we select the regularity index $s_{1}$ such that
$\max\bigg\{3\Big(\frac{1}{2}-\frac{1}{p_{1}}\Big),3\Big(\frac{1}{2}-\frac{1}{p_{2}}\Big)+1\bigg\}<s_{1}<\frac{3}{2}.$
The Sobolev embedding $H^{3(\frac{1}{2}-\frac{1}{p_{1}})}(A)\hookrightarrow L^{p_{1}}$ from Lemma 2.12 combined with the inclusion $H^{s_{1}}(A)\hookrightarrow H^{3(\frac{1}{2}-\frac{1}{p_{1}})}(A)$ (since $3(1/2-1/p_{1})<s_{1}$) yields
$\|u\|_{L^{p_{1}}}\leq C\|u\|_{H^{s_{1}}(A)}.$
For the second term, by using the Littlewood-Paley decomposition (or resolution of identity) and the Besov space embedding $H^{s_{1}}(A)=B^{s_{1}}_{2,2}(A)\hookrightarrow B^{1+3(\frac{1}{2}-\frac{1}{p_{2}})}_{p_{2},\infty}(A)$ from Lemma  2.11, followed by the triangle inequality, we obtain
$\begin{split}\|\nabla u\|_{L^{p_{2}}}\leq&\sum_{j\in\mathbb{Z}}2^{j}\|\phi_{j}(D)u\|_{L^{p_{2}}}\leq C\sum_{j\in\mathbb{Z}}2^{j+3(\frac{1}{2}-\frac{1}{p_{2}})j}\|\phi_{j}(D)u\|_{L^{2}}=C\|u\|_{B^{1+3(\frac{1}{2}-\frac{1}{p_{2}})}_{2,1}(A)}\\ \leq&C\|u\|_{H^{s_{1}}(A)}.\end{split}$
Therefore, we conclude that if $3/2<s_{2}<2$, then
$\begin{split}\|u(t)\|_{H^{s_{2}}(A)}\leq&Ct^{-\frac{s_{2}-1}{2}}\|u_{0}\|_{H^{1}(A)}+C\varepsilon^{-\frac{s_{2}}{2}}t^{1-\frac{s_{2}}{2}}\sup_{\tau\in[0,t]}\|u\|_{H^{s_{1}}(A)}^{2}<\infty.\end{split}$
Step 2. The goal of Step 2 is to prove that the time derivative of the fractional power of the solution, $\partial_{t}A^{1/2}u(t)$, belongs to $L^{2}(0,T;L^{2})$. We first prove that $\partial_{t}u\in L^{2}(0,T;L^{2})$. To this end, we start with maximal regularity estimate Proposition 2.14, which gives
$\begin{split}\|u\|_{L^{2}(0,T;H^{2}(A))}\leq&C\|u_{0}\|_{H^{1}(A)}+C\Big\|\mathbb{P}\Big((u\cdot\nabla)u\Big)\Big\|_{L^{2}(0,T;L^{2})}\\ \leq&C\|u_{0}\|_{H^{1}(A)}+CT^{\frac{1}{2}}\sup_{t\in[0,T]}\|u(t)\|_{H^{s_{1}}(A)}^{2},\quad\text{with }1<s_{1}<\frac{3}{2}.\end{split}$
Since $u\in C([0,T],H^{s_{1}}(A))$ with $s_{1}>1$, the term $\sup_{t\in[0,T]}\|u(t)\|_{H^{s_{1}}(A)}$ is finite. Consequently, this bound, combined with the maximal regularity estimate, implies that
$\partial_{t}u\in L^{2}(0,T;L^{2}).$
We now show the stronger regularity $\partial_{t}A^{\frac{1}{2}}u(t)\in L^{2}(0,T;L^{2})$. We rewrite $A^{1/2}$ on the integral equation using the identity $A^{1/2}=AB^{-\frac{1}{2}}$ (see Lemma 2.5):
$\begin{split}A^{\frac{1}{2}}u(t)=&A^{\frac{1}{2}}e^{-t\varepsilon A}u_{0}-\int_{0}^{t}Ae^{-(t-\tau)\varepsilon A}\mathbb{P}B^{-\frac{1}{2}}\Big((u\cdot\nabla)u\Big)d\tau\\ =&A^{\frac{1}{2}}e^{-t\varepsilon A}u_{0}-\varepsilon^{-1}\int_{0}^{t}(\partial_{\tau}e^{-(t-\tau)\varepsilon A})\mathbb{P}B^{-\frac{1}{2}}\Big((u\cdot\nabla)u\Big)d\tau.\end{split}$
Applying integration by parts (with respect to $\tau$ on the integral term), we obtain
$\begin{split}A^{\frac{1}{2}}u(t)=&A^{\frac{1}{2}}e^{-t\varepsilon A}u_{0}-\varepsilon^{-1}\left[e^{-(t-\tau)\varepsilon A}\mathbb{P}B^{-\frac{1}{2}}\Big((u(\tau)\cdot\nabla)u(\tau)\Big)\right]_{\tau=0}^{\tau=t}\\ &\qquad+\varepsilon^{-1}\int_{0}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}B^{-\frac{1}{2}}\partial_{\tau}\Big((u\cdot\nabla)u\Big)d\tau\\ =&A^{\frac{1}{2}}e^{-t\varepsilon A}u_{0}-\varepsilon^{-1}\mathbb{P}B^{-\frac{1}{2}}\Big((u(t)\cdot\nabla)u(t)\Big)+\varepsilon^{-1}e^{-t\varepsilon A}\mathbb{P}B^{-\frac{1}{2}}\Big((u_{0}\cdot\nabla)u_{0}\Big)\\ &\qquad+\varepsilon^{-1}\int_{0}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}B^{-\frac{1}{2}}\nabla\partial_{\tau}(u\otimes u)d\tau.\end{split}$
We now examine the time derivative $\partial_{t}A^{\frac{1}{2}}u(t)$ for each term. It is straightforward to see that the derivatives belong to $L^{2}(0,T;L^{2})$ for the first and third terms (semigroup terms), since $u_{0}$ is sufficiently regular. We directly consider the derivative of the second term:
$\partial_{t}\left[\varepsilon^{-1}\mathbb{P}B^{-\frac{1}{2}}\Big((u(t)\cdot\nabla)u(t)\Big)\right]=\varepsilon^{-1}\mathbb{P}B^{-\frac{1}{2}}\nabla\cdot\partial_{t}(u\otimes u)(t).$
Applying the boundedness of $B^{-\frac{1}{2}}\nabla$ on $L^{2}$ and the product rule, we have that
$\begin{split}\|\partial_{t}\varepsilon^{-1}\mathbb{P}B^{-\frac{1}{2}}\big((u(t)\cdot\nabla)u(t)\big)\|_{L^{2}(0,T;L^{2})}&\leq C\varepsilon^{-1}\|B^{-\frac{1}{2}}\nabla\|_{L^{2}\to L^{2}}\|\partial_{t}\big(u\otimes u\big)\|_{L^{2}(0,T;L^{2})}\\ &\leq C\varepsilon^{-1}\|u\|_{L^{\infty}(0,T;L^{\infty})}\|\partial_{t}u\|_{L^{2}(0,T;L^{2})}<\infty.\end{split}$ (3.5)
The derivative of the fourth term is given by applying the Leibniz rule for integrals:
$\begin{split}&\partial_{t}\bigg[\varepsilon^{-1}\int_{0}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}B^{-\frac{1}{2}}\nabla\partial_{\tau}(u\otimes u)~d\tau\bigg]\\ =&\varepsilon^{-1}\mathbb{P}B^{-\frac{1}{2}}\nabla\partial_{t}(u(t)\otimes u(t))-\int_{0}^{t}Ae^{-(t-\tau)\varepsilon A}\mathbb{P}B^{-\frac{1}{2}}\nabla\partial_{\tau}(u\otimes u)~d\tau.\end{split}$
The first term on the right-hand side is handled in the same manner as in (3.5). The second term on the right-hand side is bounded by maximal regularity estimate Proposition 2.14 and the previous estimate (3.5) as follows:
$\begin{split}\Big\|\varepsilon^{-1}\int_{0}^{t}Ae^{-(t-\tau)\varepsilon A}\mathbb{P}B^{-\frac{1}{2}}\nabla\partial_{\tau}(u\otimes u)~d\tau\Big\|_{L^{2}(0,T;L^{2})}\leq&C(\varepsilon)\|\mathbb{P}B^{-\frac{1}{2}}\nabla\partial_{\tau}(u\otimes u)\|_{L^{2}(0,T;L^{2})}\\ \leq&C(\varepsilon)\|u\|_{L^{\infty}(0,T;L^{\infty})}\|\partial_{t}u\|_{L^{2}(0,T;L^{2})}\\ <&\infty,\end{split}$
where $C(\varepsilon)$ is a positive constant depending on $\varepsilon$. Consequently, we conclude that $\partial_{t}A^{\frac{1}{2}}u\in L^{2}(0,T;L^{2})$.
Step 3. The goal of Step 3 is to establish the spatial regularity of the solution, specifically proving that $u$ belongs to $L^{2}(0,T;H^{3}(\Omega))$.
We start from the governing equation, the Neumann Stokes equation, written in terms of the operator $A$:
$\partial_{t}u+\varepsilon Au+\mathbb{P}\big((u\cdot\nabla)u\big)=0.$
We isolate the solution $u$ by applying the inverse operator $A^{-1}$:
$u=-A^{-1}\Big(\varepsilon^{-1}\mathbb{P}\big((u\cdot\nabla)u\big)+\varepsilon^{-1}\partial_{t}u\Big).$ (3.6)
We then apply the elliptic regularity estimate (see Lemma 4.4 in Kozono-Yanagisawa [KoYa-2009]) for the Neumann Stokes operator to obtain the spatial estimate:
$\|u\|_{H^{3}(\Omega)}\leq C\Big\|\varepsilon^{-1}\mathbb{P}((u\cdot\nabla)u)\Big\|_{H^{1}(\Omega)}+C\Big\|\varepsilon^{-1}\partial_{t}u\Big\|_{H^{1}(\Omega)}.$
Here, the Sobolev space $H^{k}(\Omega)$ is the standard Sobolev space on the domain $\Omega$, defined by a subspace of $L^{2}(\Omega)$ where the weak derivatives up to order $k$ belong to $L^{2}(\Omega)$. We utilize the boundedness of the projection $\mathbb{P}$ on $H^{1}(\Omega)$ in Lemma 2.7 and the norm equivalence between $H^{1}(\Omega)$ and $H^{1}(A)$ to estimate the nonlinear term and the time derivative term, respectively. From Step 2, we note that $A^{1/2}\partial_{t}u\in L^{2}(0,T;L^{2}(\Omega))$. By taking the $L^{2}$ norm with respect to the time variable, we obtain the following final estimate:
$\begin{split}\|u\|_{L^{2}(0,T;H^{3}(\Omega))}\leq&C\|\mathbb{P}((u\cdot\nabla)u)\|_{L^{2}(0,T;H^{1}(\Omega))}+C\|A^{1/2}\partial_{t}u\|_{L^{2}(0,T;L^{2})}\\ \leq&CT^{\frac{1}{2}}\|u\|_{L^{\infty}(0,T;H^{s}(\Omega))}\|u\|_{L^{2}(0,T;H^{2}(\Omega))}+C\|A^{\frac{1}{2}}\partial_{t}u\|_{L^{2}(0,T;L^{2})}<\infty,\end{split}$
where $3/2<s<2$. The finiteness follows from the regularity results obtained in Step 1 and Step 2.
Step 4. We claim $u\in L^{2}(0,T;H^{4}(\Omega))$. The governing equation is
$\partial_{t}u+\varepsilon Au+\mathbb{P}\big((u\cdot\nabla)u\big)=0\quad\text{ in }L^{2},\text{ for almost every }t\in(0,T).$
We apply the spectral restriction operator $\phi_{j}(D)$ to the equation:
$\partial_{t}\phi_{j}(D)u+\varepsilon A\phi_{j}(D)u+\phi_{j}(D)\mathbb{P}\big((u\cdot\nabla)u\big)=0\quad\text{ in }L^{2},\quad\text{ for almost every }t\in(0,T).$
We first justify the time differentiability of the above equation. The second term, $\varepsilon A\phi_{j}(D)u$, is time differentiable since $A\phi_{j}(D)$ is a bounded operator on $L^{2}$ and $\partial_{t}u\in L^{2}(0,T;L^{2})$. Thus $\partial_{t}(\varepsilon A\phi_{j}(D)u)\in L^{2}(0,T;L^{2})$. The regularity $\partial_{t}A^{\frac{1}{2}}u\in L^{2}(0,T;L^{2})$ from Step 2 implies that the time derivative of the nonlinear term is justified:
$\partial_{t}\left[\phi_{j}(D)\mathbb{P}\big((u\cdot\nabla)u\big)\right]=\phi_{j}(D)\mathbb{P}\big((\partial_{t}u\cdot\nabla)u\big)+\phi_{j}(D)\mathbb{P}\big((u\cdot\nabla)\partial_{t}u\big).$
Since all terms on the right-hand side of the restricted equation are time differentiable in $L^{2}(0,T;L^{2})$, we conclude that $\partial_{t}^{2}\phi_{j}(D)u$ is justified as an element of $L^{2}(0,T;L^{2})$. The twice-differentiated equation is:
$\partial_{t}^{2}\phi_{j}(D)u+\varepsilon A\partial_{t}\phi_{j}(D)u+\phi_{j}(D)\mathbb{P}\big((\partial_{t}u\cdot\nabla)u\big)+\phi_{j}(D)\mathbb{P}\big((u\cdot\nabla)\partial_{t}u\big)=0.$
This allows us to write $\partial_{t}\phi_{j}(D)u$ using the Duhamel formula. For almost every $t_{0},t$ with $0<t_{0}<t<T$, we have
$\begin{split}\phi_{j}(D)\partial_{t}u(t)=&\phi_{j}(D)e^{-(t-t_{0})\varepsilon A}\partial_{t}u(t_{0})\\ &-\phi_{j}(D)\int_{t_{0}}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}\big((\partial_{\tau}u\cdot\nabla)u+(u\cdot\nabla)\partial_{\tau}u\big)~d\tau\quad\text{ in }L^{2}.\end{split}$
Since the equality holds for all $j\in\mathbb{Z}$ and each term without $\phi_{j}(D)$ is in $L^{2}$, we may remove $\phi_{j}(D)$ to obtain an equality for $\partial_{t}u$.
$\begin{split}\partial_{t}u(t)=&e^{-(t-t_{0})\varepsilon A}\partial_{t}u(t_{0})-\int_{t_{0}}^{t}e^{-(t-\tau)\varepsilon A}\mathbb{P}\big((\partial_{\tau}u\cdot\nabla)u+(u\cdot\nabla)\partial_{\tau}u\big)~d\tau\quad\text{ in }L^{2}.\end{split}$
Furthermore, this implies that $\partial_{t}u$ is continuous with respect to $t$ in the $L^{2}$ topology.
Next, we estimate $\|A\partial_{t}u\|_{L^{2}(t_{0},T;L^{2})}$. By applying the operator $A$ and taking the $L^{2}(t_{0},T;L^{2}))$ norm on the equation for $\partial_{t}u$, we obtain the estimate:
$\begin{split}&\|A\partial_{t}u\|_{L^{2}(t_{0},T;L^{2})}\\ \leq&C\|\partial_{t}u(t_{0})\|_{H^{1}(A)}+C\left\|\int_{t_{0}}^{t}Ae^{-(t-\tau)\varepsilon A}\mathbb{P}\big((\partial_{\tau}u\cdot\nabla)u+(u\cdot\nabla)\partial_{\tau}u\big)~d\tau\right\|_{L^{2}(t_{0},T;L^{2})}\\ \leq&C\|\partial_{t}u(t_{0})\|_{H^{1}(A)}+C\|(\partial_{\tau}u\cdot\nabla)u+(u\cdot\nabla)\partial_{\tau}u\|_{L^{2}(t_{0},T;L^{2})}\\ \leq&C\|\partial_{t}u(t_{0})\|_{H^{1}(A)}+C\|\partial_{\tau}u\|_{L^{2}(t_{0},T;L^{6})}\|\nabla u\|_{L^{\infty}(t_{0},T;L^{3})}+C\|u\|_{L^{\infty}(t_{0},T;L^{\infty})}\|\nabla\partial_{\tau}u\|_{L^{2}(t_{0},T;L^{2})},\end{split}$
where we have used maximal regularity estimate Proposition 2.14. We choose $t_{0}\in(0,T)$ such that $A^{\frac{1}{2}}\partial_{t}u(t_{0})\in L^{2}$, which is possible since $\partial_{t}u\in L^{2}(0,T;H^{1}(A))$ by Step 2, and the first term $\|\partial_{t}u(t_{0})\|_{H^{1}(A)}$ is finite. For the second term, by Sobolev embedding $H^{1}(\Omega)\hookrightarrow L^{6}$ and $A^{\frac{1}{2}}\partial_{t}u\in L^{2}(0,T;L^{2})$ from Step 2, we have
$\|\partial_{t}u\|_{L^{2}(t_{0},T;L^{6})}\leq C\|\partial_{t}u\|_{L^{2}(t_{0},T;H^{1}(\Omega))}\leq C\|\partial_{t}u\|_{L^{2}(t_{0},T;H^{1}(A))}<\infty.$
We also estimate $\|\nabla u\|_{L^{3}}$ using the embedding theorem Lemma 2.12:
$\|\nabla u\|_{L^{3}}\leq C\|u\|_{H^{s}(A)},\quad\frac{3}{2}<s<2,$
which is finite by Step 1. For the fourth term, the estimate is:
$\|u\|_{L^{\infty}(t_{0},T;L^{\infty})}\|\nabla\partial_{\tau}u\|_{L^{2}(t_{0},T;L^{2})}\leq C\|u\|_{L^{\infty}(t_{0},T;H^{s}(A))}\|A^{\frac{1}{2}}\partial_{\tau}u\|_{L^{2}(t_{0},T;L^{2})}<\infty.$
Therefore, we conclude that $A\partial_{t}u\in L^{2}(t_{0},T;L^{2})$.
Finally, we estimate $\|u\|_{L^{2}(t_{0},T;H^{4}(\Omega))}$. It follows from the elliptic estimate in Lemma 2.6 and the equations that
$\|u\|_{H^{4}(\Omega)}=\|A^{-1}Au\|_{H^{4}(\Omega)}\leq C\|Au\|_{H^{2}(\Omega)}\leq C\|\mathbb{P}((u\cdot\nabla)u)\|_{H^{2}(\Omega)}+C\|\partial_{t}u\|_{H^{2}(\Omega)}.$
The $H^{4}(\Omega)$ norm of the solution is then estimated as:
$\begin{split}\|u\|_{H^{4}(\Omega)}\leq&C(\|u\|_{L^{6}}+\|\nabla^{2}u\|_{L^{6}})\|\nabla u\|_{L^{3}}+C\|u\|_{L^{\infty}}\|u\|_{H^{3}(\Omega)}+C\|A\partial_{t}u\|_{L^{2}}\\ \leq&C\|u\|_{H^{3}(\Omega)}\|u\|_{H^{s}(A)}+C\|A\partial_{t}u\|_{L^{2}},\quad\text{ with }\frac{3}{2}<s<2.\end{split}$
By taking the $L^{2}$ norm in $t$, we obtain:
$\|u\|_{L^{2}(t_{0},T;H^{4}(\Omega))}\leq C\|u\|_{L^{\infty}(t_{0},T;H^{s}(A))}\|u\|_{L^{2}(t_{0},T;H^{3}(\Omega))}+C\|A\partial_{t}u\|_{L^{2}(t_{0},T;L^{2})}<\infty.$
The finiteness follows from the estimates in Step 1, Step 3, and the conclusion $A\partial_{t}u\in L^{2}(t_{0},T;L^{2})$ just established. $\Box$
4. Uniform Boundedness with respect to $\varepsilon$
The goal of this section is to establish the uniform boundedness of the solution $u_{\varepsilon}$ in $L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$ with respect to $\varepsilon$. To this end, we provide two propositions: Proposition 4.1 ensures the regularity in $L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$, while Proposition 4.2 establishes the uniform boundedness and the time continuity.
Proposition 4.1.
Let $0<\varepsilon<1$ and $u_{0}\in B^{\frac{5}{2}}_{2,1}(A)$. Then, there exists a time $T>0$ such that the integral equation (3.1) admits a unique solution $u_{\varepsilon}$ satisfying
$u_{\varepsilon}\in C([0,T],H^{1}(A))\cap L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A)).$
Proof. Step 1. It follows from Propositions 3.1, 3.2 that a unique solution $u_{\varepsilon}\in C([0,T],H^{1}(A))$ exists and possesses the following regularity properties:
$\begin{split}&u_{\varepsilon}\in L^{2}((0,T],H^{2}(A))\cap L^{2}((0,T],H^{3}(\Omega))\cap L^{2}_{\rm loc}((0,T],H^{4}(\Omega)),\quad\\ &\partial_{t}u_{\varepsilon}\in L^{2}_{\rm loc}((0,T],H^{2}(A)).\end{split}$
Applying the Laplacian $\Delta$ to the partial differential equations yields the following differentiated equations:
$\partial_{t}\Delta u_{\varepsilon}-\varepsilon\Delta^{2}u_{\varepsilon}+(u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}+N(u_{\varepsilon},u_{\varepsilon})=0\quad\text{ in }L^{2}(0,T;L^{2}),$
and (2.2) in the $L^{2}$ setting yields that the nonlinear remainder term $N(u,v)$ is given by
$N(u,v):=(\Delta u\cdot\nabla)v+2(\nabla u\cdot\nabla)\nabla v+\nabla{\rm div}\,\big((u\cdot\nabla)v\big).$
We note that the term $2(\nabla u\cdot\nabla)\nabla v$ is understood componentwise. Specifically, for the $k$-th component ($k=1,2,3$), it is defined by:
$\Big(2(\nabla u\cdot\nabla)\nabla v\Big)_{k}=\sum_{i,j=1}^{3}2(\partial_{x_{i}}u_{j})(\partial_{x_{j}}\partial_{x_{i}}v_{k}).$
We now apply the spectral restriction operator $\psi_{j}(D)$ to the differentiated equation to obtain:
$\psi_{j}(D)\Big(\partial_{t}\Delta u_{\varepsilon}-\varepsilon\Delta^{2}u_{\varepsilon}+(u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}+N(u_{\varepsilon},u_{\varepsilon})\Big)=0\quad\text{ in }L^{2}(0,T;L^{2}).$ (4.1)
In the subsequent steps, we will multiply the equation (4.1) by an appropriate test function, such as $\Delta u_{\varepsilon}$, and integrate in space to obtain energy estimates. We structure the proof as follows: we analyze the terms involving $\partial_{t}\Delta u_{\varepsilon}$, $\varepsilon\Delta^{2}u_{\varepsilon}$, and $(u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}$ in Step 2; we estimate the nonlinear remainder term $N(u_{\varepsilon},u_{\varepsilon})$ in Step 3; and finally we prove the existence of $T$ such that $u\in L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$ in Step 4.
Step 2. We analyze the terms in the equation (4.1) multiplied by the test function $\Delta u_{\varepsilon}$. The time derivative term is converted into a total time derivative:
$\begin{split}\int_{\Omega}\psi_{j}(D)\partial_{t}\Delta u_{\varepsilon}\cdot\Delta u_{\varepsilon}\,dx&=\int_{\Omega}\partial_{t}\left(\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\right)\cdot\left(\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\right)\,dx\\ &=\dfrac{1}{2}\partial_{t}\int_{\Omega}|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}|^{2}\,dx=\dfrac{1}{2}\partial_{t}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}^{2}.\end{split}$
For the viscous term, applying Lemma 2.15, we ensure a favorable sign:
$\int_{\Omega}\Big(\psi_{j}(D)\big(\varepsilon\Delta^{2}u_{\varepsilon}\big)\Big)\cdot\Delta u_{\varepsilon}\,dx=\varepsilon\int_{\Omega}\Big(\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\Big)\cdot\Big(\psi_{j}(D)^{\frac{1}{2}}\Delta^{2}u_{\varepsilon}\Big)\,dx\geq 0.$
We estimate the convection term. Using the symmetry of the operator $\psi_{j}(D)$, we first shift the operator:
$\int_{\Omega}\Big(\psi_{j}(D)\big((u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}\big)\Big)\cdot\Delta u_{\varepsilon}\,dx=\int_{\Omega}\big((u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}\big)\cdot\psi_{j}(D)\Delta u_{\varepsilon}\,dx.$
The next step involves an identity based on integration by parts together with the boundary condition $u_{\varepsilon}\cdot\nu=0$:
$\int_{\Omega}\big((u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}\big)\cdot\psi_{j}(D)\Delta u_{\varepsilon}\,dx=-\int_{\Omega}\big((u_{\varepsilon}\cdot\nabla)\psi_{j}(D)\Delta u_{\varepsilon}\big)\cdot\Delta u_{\varepsilon}\,dx.$
Then, using the specific definition of the operator $\psi_{j}(D)$ via the spectral operator $B$:
$\psi_{j}(D)=\sum_{l=0}^{1}(-1)^{l+1}(1+2^{-2j-2l}B)^{-1},$
we write
$\int_{\Omega}\Big(\psi_{j}(D)\big((u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}\big)\Big)\cdot\Delta u_{\varepsilon}\,dx=-\sum_{l=0}^{1}(-1)^{l+1}\int_{\Omega}\big((u_{\varepsilon}\cdot\nabla)(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}\big)\cdot\Delta u_{\varepsilon}\,dx.$
We use the identity $\Delta=\nabla\text{div}-\text{rot}\,\text{rot}$ and the divergence-free condition of the term $(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}$ to write:
$\begin{split}\Delta u_{\varepsilon}=&(1+2^{-2j-2l}B)(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}\\ =&(1+2^{-2j-2l}\text{rot}\,\text{rot})(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}.\end{split}$
Substituting this back and applying integration by parts:
$\begin{split}&\int_{\Omega}\Big(\psi_{j}(D)\big((u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}\big)\Big)\cdot\Delta u_{\varepsilon}\,dx\\ &=\sum_{l=0}^{1}(-1)^{l+1}2^{-2j-2l}\int_{\Omega}\text{rot}\Big((u_{\varepsilon}\cdot\nabla)(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}\Big)\cdot\text{rot}\big((1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}\big)\,dx.\end{split}$
Integration by parts and the boundary condition $u_{\varepsilon}\cdot\nu=0$ imply that
$\int_{\Omega}\big((u_{\varepsilon}\cdot\nabla)U\big)\cdot U\,dx=0,\quad\text{ with }U={\rm rot}\big((1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}\big),$
and we obtain:
$\begin{split}&\int_{\Omega}\Big(\psi_{j}(D)\big((u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}\big)\Big)\cdot\Delta u_{\varepsilon}\,dx\\ &=\sum_{l=0}^{1}(-1)^{l+1}2^{-2j-2l}\int_{\Omega}\Big\{\big((\widetilde{\rm rot}\,u_{\varepsilon}\cdot\nabla)(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}\big)\Big\}\cdot U\,dx,\end{split}$
where $U={\rm rot}(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}$ and the notation $\widetilde{\rm rot}\,u$ is defined as the remainder of the vector calculus identity:
$(\widetilde{\rm rot}\,u_{\varepsilon}\cdot\nabla)v:={\rm rot}\big((u_{\varepsilon}\cdot\nabla)v\big)-(u_{\varepsilon}\cdot\nabla){\rm rot}\,v.$
Applying the Hölder inequality to the resulting term:
$\begin{split}&\Big|\int_{\Omega}\Big(\psi_{j}(D)\big((u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}\big)\Big)\cdot\Delta u_{\varepsilon}\,dx\Big|\\ &\leq\sum_{l=0}^{1}C2^{-2j-2l}\|\widetilde{\rm rot}\,u_{\varepsilon}\|_{L^{\infty}}\|\nabla(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}\|_{L^{2}}\|U\|_{L^{2}}\\ &\leq C2^{-2j}\|\nabla u_{\varepsilon}\|_{L^{\infty}}\cdot\|\nabla(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}\|_{L^{2}}\cdot\|U\|_{L^{2}}.\end{split}$
It follows from (2.6) that
$\begin{split}\|\nabla(1+2^{-2j-2l}B)^{-1}\Delta u_{\varepsilon}\|_{L^{2}},\|U\|_{L^{2}}\leq C2^{j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}},\end{split}$
which implies that
$\Big|\int_{\Omega}\Big(\psi_{j}(D)\big((u_{\varepsilon}\cdot\nabla)\Delta u_{\varepsilon}\big)\Big)\cdot\Delta u_{\varepsilon}\,dx\Big|\leq C\|\nabla u_{\varepsilon}\|_{L^{\infty}}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}^{2}.$
Step 3. We estimate the integral involving the nonlinear remainder term $N(u_{\varepsilon},u_{\varepsilon})$:
$\int_{\Omega}\Big(\psi_{j}(D)N(u_{\varepsilon},u_{\varepsilon})\Big)\cdot\Delta u_{\varepsilon}\,dx.$
Due to the orthogonality property (see Lemma 2.1) of the pressure term $\nabla{\rm div}\,\big((u_{\varepsilon}\cdot\nabla)u_{\varepsilon}\big)$ with a divergence-free vector $\psi_{j}(D)\Delta u_{\varepsilon}$, the integral simplifies to:
$\int_{\Omega}\Big(\psi_{j}(D)N(u_{\varepsilon},u_{\varepsilon})\Big)\cdot\Delta u_{\varepsilon}\,dx=\int_{\Omega}\Big\{(\Delta u_{\varepsilon}\cdot\nabla)u_{\varepsilon}+2(\nabla u_{\varepsilon}\cdot\nabla)\nabla u_{\varepsilon}\Big\}\cdot\psi_{j}(D)\Delta u_{\varepsilon}\,dx.$
We focus on estimating the term involving $(\Delta u_{\varepsilon}\cdot\nabla)u_{\varepsilon}$, as the second term $2(\nabla u_{\varepsilon}\cdot\nabla)\nabla u_{\varepsilon}$ can be treated analogously using similar product decomposition techniques.
We use the low-frequency cut-off operator $S_{j}=\sum_{k\leq j}\phi_{k}(D)$ to decompose the term $(\Delta u_{\varepsilon}\cdot\nabla)u_{\varepsilon}$:
$\begin{split}(\Delta u_{\varepsilon}\cdot\nabla)u_{\varepsilon}&=(\Delta u_{\varepsilon}\cdot\nabla)(1-S_{j})u_{\varepsilon}+((1-S_{j})\Delta u_{\varepsilon}\cdot\nabla)S_{j}u_{\varepsilon}+(S_{j}\Delta u_{\varepsilon}\cdot\nabla)S_{j}u_{\varepsilon}\\ &=N_{1}+N_{2}+N_{3}.\end{split}$
Next, we estimate the nonlinear terms. Applying $L^{3}\times L^{6}\times L^{2}$ estimate, $H^{1}(\Omega)\hookrightarrow L^{3}$, $H^{1}(\Omega)\hookrightarrow L^{6}$, and (2.5), the term $N_{1}$ is estimated as:
$\begin{split}\Big|\int_{\Omega}\Big(\psi_{j}(D)N_{1}\Big)\cdot\Delta u_{\varepsilon}~dx\Big|\leq&C\|\Delta u_{\varepsilon}\|_{L^{3}}\|\nabla(1-S_{j})u_{\varepsilon}\|_{L^{6}}\|\psi_{j}(D)\Delta u_{\varepsilon}\|_{L^{2}}\\ \leq&C\|u_{\varepsilon}\|_{H^{3}(\Omega)}\|\nabla(1-S_{j})u_{\varepsilon}\|_{H^{1}(\Omega)}\|\psi_{j}(D)\Delta u_{\varepsilon}\|_{L^{2}}\\ \leq&C\|u_{\varepsilon}\|_{H^{3}(\Omega)}\|(1-S_{j})u_{\varepsilon}\|_{H^{2}(A)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}.\end{split}$ (4.2)
For $N_{2}$, we use $L^{2}\times L^{\infty}\times L^{2}$ esimate, Lemma 2.11 and (2.5):
$\begin{split}&\Big|\int_{\Omega}\Big(\psi_{j}(D)N_{2}\Big)\cdot\Delta u_{\varepsilon}~dx\Big|\\ \leq&C\|(1-S_{j})\Delta u_{\varepsilon}\|_{L^{2}}\|\nabla S_{j}u_{\varepsilon}\|_{L^{\infty}}\|\psi_{j}(D)\Delta u_{\varepsilon}\|_{L^{2}}\\ \leq&C\|(1-S_{j})u_{\varepsilon}\|_{H^{2}(A)}\sum_{k\leq j}2^{\frac{5}{2}j}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}.\end{split}$ (4.3)
For the term $N_{3}=(S_{j}\Delta u_{\varepsilon}\cdot\nabla)S_{j}u_{\varepsilon}$, we use the vector calculus identity and boundary conditions. Since $\psi_{j}(D)\Delta u_{\varepsilon}=\Delta\psi_{j}(D)u_{\varepsilon}$ and $\nabla\cdot(\psi_{j}(D)u_{\varepsilon})=0$, we have
$\psi_{j}(D)\Delta u_{\varepsilon}=\text{rot}\,\text{rot}\,\psi_{j}(D)u_{\varepsilon}.$
We apply integration by parts using the curl operator and ${\rm rot\,}\psi_{j}(D)u_{\varepsilon}\times\nu=0$ on the boundary:
$\int_{\Omega}\Big(\psi_{j}(D)N_{3}\Big)\cdot\Delta u_{\varepsilon}\,dx=\int_{\Omega}\text{rot}\,N_{3}\cdot\text{rot}\,\psi_{j}(D)u_{\varepsilon}\,dx.$
By $L^{2}\times L^{\infty}\times L^{2}$ estimate:
$\begin{split}&\Big|\int_{\Omega}\Big(\psi_{j}(D)N_{3}\Big)\cdot\Delta u_{\varepsilon}~dx\Big|\\ \leq&\big(\|\nabla S_{j}\Delta u_{\varepsilon}\|_{L^{2}}\|\nabla S_{j}u_{\varepsilon}\|_{L^{\infty}}+\|S_{j}\Delta u_{\varepsilon}\|_{L^{6}}\|\nabla^{2}S_{j}u_{\varepsilon}\|_{L^{3}}\big)\|{\rm rot\,}\psi_{j}(D)u_{\varepsilon}\|_{L^{2}}.\end{split}$
On the term $\|{\rm rot\,}\psi_{j}(D)u_{\varepsilon}\|_{L^{2}}$, we estimate
$\|{\rm rot\,}\psi_{j}(D)u_{\varepsilon}\|_{L^{2}}\leq C\|B^{\frac{1}{2}}\psi_{j}(D)u_{\varepsilon}\|_{L^{2}}.$
We notice that
$\begin{split}B^{\frac{1}{2}}\psi_{j}(D)u_{\varepsilon}=&\dfrac{3}{4}\cdot 2^{-2j}B^{\frac{1}{2}}(1+2^{-2j-2}B)^{-1}(1+2^{-2j}B)^{-1}Bu_{\varepsilon}\\ =&\dfrac{\sqrt{3}}{2}\cdot 2^{-j}(1+2^{-2j-2}B)^{-\frac{1}{2}}(1+2^{-2j}B)^{-\frac{1}{2}}\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon},\end{split}$
and that $(1+2^{-2j-2}B)^{-\frac{1}{2}}(1+2^{-2j}B)^{-\frac{1}{2}}$ is a bounded operator uniform on $L^{2}$ with respect to $j$, which yield that
$\|{\rm rot\,}\psi_{j}(D)u\|_{L^{2}}\leq C2^{-j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}},$
Substituting this back and applying Lemma 2.11, we obtain the final estimate for $N_{3}$:
$\begin{split}&\Big|\int_{\Omega}\Big(\psi_{j}(D)N_{3}\Big)\cdot\Delta u_{\varepsilon}~dx\Big|\\ \leq&C\Big(\sum_{k\leq j}2^{3k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}\Big)\Big(\sum_{l\leq j}2^{\frac{5}{2}l}\|\phi_{l}(D)u_{\varepsilon}\|_{L^{2}}\Big)\cdot 2^{-j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}.\end{split}$ (4.4)
Step 4. By Steps 1, 2 and the Sobolev embedding $H^{2}(\Omega)\hookrightarrow L^{\infty}$, we obtain:
$\begin{split}\partial_{t}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}^{2}\leq&C\|\nabla u_{\varepsilon}\|_{L^{\infty}}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}^{2}+\Big|\int_{\Omega}\Big(\psi_{j}(D)N(u_{\varepsilon},u_{\varepsilon})\Big)\cdot\Delta u_{\varepsilon}\Big|\\ \leq&C\|u_{\varepsilon}\|_{H^{3}(\Omega)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}^{2}+\Big|\int_{\Omega}\Big(\psi_{j}(D)N(u_{\varepsilon},u_{\varepsilon})\Big)\cdot\Delta u_{\varepsilon}\Big|.\end{split}$ (4.5)
Assuming $\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}>0$, we divide the above by this norm.
Now, fix a natural number $J$ and consider the partial sum over $j\leq J$. Dividing (4.5) by $\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}$, multiplying by $2^{\frac{1}{2}j}$, integrating in time, and summing over $j\leq J$, the time derivative term becomes:
$\sum_{j\leq J}2^{\frac{1}{2}j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}(t)\|_{L^{2}}-\sum_{j\leq J}2^{\frac{1}{2}j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}(0)\|_{L^{2}}.$
For the first term of the right-hand-side of (4.5), we have:
$\begin{split}&\sum_{j\leq J}2^{\frac{1}{2}}\int_{0}^{t}\|u_{\varepsilon}\|_{H^{3}(\Omega)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}~d\tau\\ \leq&CT^{\frac{1}{2}}\|u_{\varepsilon}\|_{L^{2}(0,T;H^{3}(\Omega))}\sup_{t\in[0,T]}\sum_{j\leq J}2^{\frac{1}{2}j}\|\psi_{j}(D)\Delta u_{\varepsilon}(t)\|_{L^{2}}.\end{split}$
Applying (4.2), it follows that:
[…]
Similarly, for (4.3) and (4.4)
[…]
and
[…]
We introduce the following quantities, which arise from the difference between $\phi_{j}$ and $\psi_{j}$:
$A_{J}(t):=\sup_{\tau\in(0,t)}\sum_{j\leq J}2^{\frac{5}{2}j}\|\phi_{j}(D)u_{\varepsilon}(\tau)\|_{L^{2}},\quad\widetilde{A}_{J}(t):=\sup_{\tau\in(0,t)}\sum_{j\leq J}2^{\frac{5}{2}j}\|\psi_{j}(D)u_{\varepsilon}(\tau)\|_{L^{2}}.$
Note that $u_{\varepsilon}\in L^{2}(0,T;H^{3}(\Omega))$, $\sup_{J}\|(1-S_{J})u_{\varepsilon}\|_{L^{2}(0,T;H^{2}(A))}<\infty$, and $A_{J}(t)\leq C\widetilde{A}_{J}(t)$. Applying the estimates above to (4.5) implies the existence of a constant $C_{0}$, depending on $u_{0}$ and $\varepsilon$, such that:
$\begin{split}\widetilde{A}_{J}(t)\leq&\widetilde{A}_{J}(0)+C_{0}+C_{0}(t^{\frac{1}{2}}+t)(A_{J}(t)+A_{J}(t)^{2})\\ \leq&\widetilde{A}_{J}(0)+C_{0}+C_{0}(t^{\frac{1}{2}}+t)(\widetilde{A}_{J}(t)+\widetilde{A}_{J}(t)^{2}).\end{split}$
for all $J$. A standard continuity argument then ensures that there exists $T_{0}>0$ such that
$\widetilde{A}_{J}(t)\leq 2(\widetilde{A}_{J}(0)+C_{0})\quad\text{ for all }J\text{ and }t\in[0,T_{0}].$
Taking the limit $J\to\infty$, the equivalence of Besov norms from Lemma 2.8 yields:
$\|u_{\varepsilon}(t)\|_{B^{\frac{5}{2}}_{2,1}(A)}\leq 2(\widetilde{A}_{J}(0)+C_{0})\quad\text{ for all }t\in[0,T_{0}].$
We then obtain the regularity $u_{\varepsilon}$ in $L^{\infty}(0,T_{0};B^{\frac{5}{2}}_{2,1}(A))$. $\Box$
Proposition 4.2.
Let $0<\varepsilon<1$ and $u_{0}\in B^{\frac{5}{2}}_{2,1}(A)$. Then there exists an existence time, $T$, of the solution $u_{\varepsilon}\in C([0,T],H^{1}(A))$ of (3.1), which depends on $\|u_{0}\|_{B^{\frac{5}{2}}_{2,1}(A)}$ and is independent of $\varepsilon$, such that the solution $u_{\varepsilon}$ satisfies the following two properties:
  1. (1)
    Uniform boundedness: $\sup_{0<\varepsilon<1}\sum_{j\in\mathbb{Z}}2^{\frac{5}{2}j}\|\phi_{j}(D)u_{\varepsilon}(t)\|_{L^{\infty}(0,T;L^{2})}<\infty,$
  2. (2)
    Regularity: $u_{\varepsilon}\in C([0,T],B^{\frac{5}{2}}_{2,1}(A))$.
Proof. Step 1. The existence of a unique solution $u_{\varepsilon}\in C([0,T],H^{1}(A))\cap L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$ is already established in Proposition 4.1, where $T$ depends on $\varepsilon$. Starting from (4.5), we derive several inequalities in this step modify the argument used in the proof of Proposition 4.1.
The first inequality of (4.5) combined with Lemma 2.11 yields:
$\partial_{t}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}^{2}\leq C\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}^{2}+\Big|\int_{\Omega}\Big(\psi_{j}(D)N(u_{\varepsilon},u_{\varepsilon})\Big)\cdot\Delta u_{\varepsilon}\Big|.$
We estimate the integral involving the nonlinear remainder term $N(u_{\varepsilon},u_{\varepsilon})$. For $N_{1}$, we modify (4.2) by using the inequality $\|\Delta u_{\varepsilon}\|_{L^{3}}\leq C\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}$, ensured by Lemma 2.11. We then have
$\begin{split}\Big|\int_{\Omega}\Big(\psi_{j}(D)N_{1}\Big)\cdot\Delta u_{\varepsilon}\,dx\Big|&\leq\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}}\|(1-S_{j})u_{\varepsilon}\|_{H^{2}(A)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}\\ &\leq C\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}}\sum_{k>j}2^{2k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}.\end{split}$
Similarly, for (4.3) and (4.4)
$\begin{split}\Big|\int_{\Omega}\Big(\psi_{j}(D)N_{2}\Big)\cdot\Delta u_{\varepsilon}\,dx\Big|\leq&C\left(\sum_{k>j}2^{2k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}\right)\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}\end{split}$
and
$\begin{split}\Big|\int_{\Omega}\Big(\psi_{j}(D)N_{3}\Big)\cdot\Delta u_{\varepsilon}\,dx\Big|&\leq C2^{-j}\left(\sum_{k\leq j}2^{3k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}\right)\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}.\end{split}$
The analogous term $2(\nabla u_{\varepsilon}\cdot\nabla)\nabla u_{\varepsilon}$ is treated similarly and yields a comparable estimate.
Step 2. By combining the estimates obtained in the previous step, the energy inequality for the localized equation is given by:
$\begin{split}\dfrac{1}{2}\partial_{t}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}^{2}\leq&C\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}^{2}\\ &+C\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\sum_{k>j}2^{2k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}\\ &+C2^{-j}\Big(\sum_{k\leq j}2^{3k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}\Big)\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}.\end{split}$
Dividing the inequality by $\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}$, we obtain the differential inequality for the $L^{2}$-norm:
$\begin{split}\partial_{t}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}\leq&C\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}+C\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\sum_{k>j}2^{2k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}\\ &+C2^{-j}\Big(\sum_{k\leq j}2^{3k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}\Big)\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}.\end{split}$ (4.6)
By integrating this inequality in time from 0 to $t$, multiplying by $2^{\frac{1}{2}j}$, and then summing over all $j\in\mathbb{Z}$, we arrive at the following estimate:
$\begin{split}\sum_{j\in\mathbb{Z}}2^{\frac{1}{2}j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}(t)\|_{L^{2}}\leq&\sum_{j\in\mathbb{Z}}2^{\frac{1}{2}j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}(0)\|_{L^{2}}\\ &+C\int_{0}^{t}\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\sum_{j\in\mathbb{Z}}2^{\frac{1}{2}j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}\,d\tau\\ &+C\int_{0}^{t}\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\sum_{k\in\mathbb{Z}}2^{\frac{5}{2}k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}d\tau.\end{split}$
Recognizing that the summations above give the $B^{\frac{5}{2}}_{2,1}(A)$ norm (see Lemma 2.8), the estimate simplifies to:
$\|u_{\varepsilon}(t)\|_{B^{\frac{5}{2}}_{2,1}(A)}\leq C\|u_{0}\|_{B^{\frac{5}{2}}_{2,1}(A)}+CT\|u_{\varepsilon}\|_{L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))}^{2},$ (4.7)
where the constant $C$ is independent of $\varepsilon$.
Moreover, we discuss the independency of the existence time $T$ with respect to $\varepsilon$. We know from the proof of Proposition 3.1 that we can construct solutions as long as the $H^{1}(A)$ norm is bounded. The estimate (4.7), coupled with $B^{\frac{5}{2}}_{2,1}(A)\hookrightarrow H^{1}(A)$ and a standard continuity argument implies that the existence time $T$ can be chosen independent of $\varepsilon$, only depending on the initial data:
$T\geq\frac{C}{\|u_{0}\|_{B^{\frac{5}{2}}_{2,1}(A)}},$
with $C$ independent of $\varepsilon$.
Step 3. We prove $u\in C([0,T],B^{\frac{5}{2}}_{2,1}(A))$. Integrating (4.6), taking the supremum with respect to the time variable, and summing over $j\in\mathbb{Z}$:
$\begin{split}\sum_{j\in\mathbb{Z}}2^{\frac{1}{2}j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}(t)\|_{L^{\infty}(0,T;L^{2})}\leq&C\|u_{0}\|_{B^{\frac{5}{2}}_{2,1}(A)}\\ &+C\int_{0}^{T}\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\sum_{j\in\mathbb{Z}}2^{\frac{1}{2}j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u_{\varepsilon}\|_{L^{2}}\,d\tau\\ &+C\int_{0}^{T}\|u_{\varepsilon}\|_{B^{\frac{5}{2}}_{2,1}(A)}\sum_{k\in\mathbb{Z}}2^{\frac{5}{2}k}\|\phi_{k}(D)u_{\varepsilon}\|_{L^{2}}d\tau\\ \leq&C\|u_{0}\|_{B^{\frac{5}{2}}_{2,1}(A)}+CT\|u_{\varepsilon}\|_{L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))}^{2}<\infty.\end{split}$
We know $u\in C([0,T],H^{1}(A))$, and Lemma A.2 implies $u\in C([0,T],B^{\frac{5}{2}}_{2,1}(A))$. $\Box$
5. Proof of Theorem 1.1
Let $\varepsilon>0$ and $u_{0}\in B^{\frac{5}{2}}_{2,1}(A)$. The existence of a solution of the integral equation (3.1) is ensured by Proposition 3.1, and its regularity follows from Propositions 3.2 and 4.2. Specifically, there exists a unique solution $u_{\varepsilon}$ satisfying
$\begin{split}&u_{\varepsilon}\in C([0,T],B^{\frac{5}{2}}_{2,1}(A))\cap C^{1}([0,T],H^{1}(A))\cap L^{2}(0,T;H^{3}(\Omega)),\\ &\partial_{t}u_{\varepsilon}\in L^{2}_{\rm loc}((0,T],H^{2}(A)),\end{split}$
and the equation
$\partial_{t}u_{\varepsilon}-\varepsilon\Delta u_{\varepsilon}+\mathbb{P}\Big((u_{\varepsilon}\cdot\nabla)u_{\varepsilon}\Big)=0.$
Since the solution $u_{\varepsilon}$ satisfies a sufficient regularity, for any test function $\varphi\in C([0,T],H^{2}(A))\cap C^{1}((0,T),L^{2}_{\sigma})$ with ${\rm supp\,}\varphi\subset(0,T)\times\overline{\Omega}$, the following integral identity holds:
$\int_{0}^{T}\int_{\Omega}\Big\{-u_{\varepsilon}\cdot\partial_{t}\varphi+\varepsilon u_{\varepsilon}\cdot\Delta\varphi-\Big(\big(u_{\varepsilon}\cdot\nabla\big)\varphi\Big)\cdot u_{\varepsilon}\Big\}\,dxdt=0.$ (5.1)
Compactness property in Lemma A.1 and the uniform boundedness established in Proposition 4.2 imply that, as $\varepsilon\to 0$, a subsequence converges to a limit $u$ in the weak-* topology of $\widetilde{L}^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$, where $\widetilde{L}^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$ is a Chemin-Lernar space (see [BaChDa_2011]) equipped with the norm defined as follows.
$\|u_{\varepsilon}\|_{\widetilde{L}^{\infty}(0,TB^{\frac{5}{2}}_{2,1}(A))}:=\sum_{j\in\mathbb{Z}}2^{\frac{5}{2}j}\|\phi_{j}(D)u_{\varepsilon}\|_{L^{\infty}(0,T;L^{2})}.$
Noting the equivalence of the norm with replacing $\phi_{j}$ by $\psi_{j}$ analogously to Lemma 2.8, we get the uniform bounds of the spacial derivative.
$\sup_{0<\varepsilon<1}\|\nabla u_{\varepsilon}\|_{L^{\infty}(0,T;L^{2})}<\infty.$
We also observe the uniform boundedness of the time derivative with respect to $0<\varepsilon<1$. Indeed, the uniformity follows from
$\begin{split}\|\partial_{t}u_{\varepsilon}\|_{L^{\infty}(0,T;L^{2})}&\leq\varepsilon\|\Delta u_{\varepsilon}\|_{L^{\infty}(0,T;L^{2})}+\big\|\mathbb{P}\big((u_{\varepsilon}\cdot\nabla)u_{\varepsilon}\big)\big\|_{L^{\infty}(0,T;L^{2})}\\ &\leq C\varepsilon\|u_{\varepsilon}\|_{L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))}+C\|u_{\varepsilon}\|_{L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))}^{2}.\end{split}$
The uniform boundedness of both time and spatial derivatives enables us to extract a subsequence that converges to $u$ in $L^{2}((0,T)\times\Omega)$. Passing to the limit in (5.1), we obtain
$\int_{0}^{T}\int_{\Omega}\Big\{-u\cdot\partial_{t}\varphi-\Big(\big(u\cdot\nabla\big)\varphi\Big)\cdot u\Big\}\,dxdt=0,$
which implies
$\int_{0}^{T}\int_{\Omega}\Big\{-u\cdot\partial_{t}\varphi+\Big(\big(u\cdot\nabla\big)u\Big)\cdot\varphi\Big\}\,dxdt=0.$
The regularity $u\in\widetilde{L}^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))\subset L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$ ensures that the nonlinear term $(u\cdot\nabla)u$ belongs to $L^{\infty}(0,T;W^{1,3}(\Omega))$. Consequently, the time derivative $\partial_{t}u$ exists as an element of $L^{\infty}(0,T;W^{1,3}(\Omega))$, and we have
$\partial_{t}u+\mathbb{P}\big((u\cdot\nabla)u\big)=0\quad\text{ in }W^{1,3}(\Omega)\text{ for almost every }t\in(0,T).$
Next, we prove the uniqueness of the solution in $L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$. Let $u$ and $v$ be two solutions with the same initial data. Following the same justification for the time derivatives as above, we consider the energy estimate for the difference uvu-v:
$\frac{1}{2}\partial_{t}\|u-v\|_{L^{2}}^{2}=\left|\int_{\Omega}\Big(\big((u-v)\cdot\nabla\big)u\Big)\cdot(u-v)\,dx\right|\leq\|\nabla u\|_{L^{\infty}}\|u-v\|_{L^{2}}^{2}.$
Since $\|\nabla u\|_{L^{\infty}}\leq C\|u\|_{B^{\frac{5}{2}}_{2,1}(A)}$ and $u\in L^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$, the Gronwall inequality yields uv=0u-v=0.
It remains to establish the time continuity of the solution. By Lemma A.1, we can show that $u$ satisfies
$\sum_{j\in\mathbb{Z}}2^{\frac{1}{2}j}\|\psi_{j}(D)^{\frac{1}{2}}\Delta u\|_{L^{\infty}(0,T;L^{2})}<\infty.$
Combining this with $\partial_{t}u\in L^{\infty}(0,T;W^{1,3}(\Omega))$ obtained before, we have $u\in C([0,T],L^{2})$. Finally, Lemma A.2 ensures the time continuity in $B^{\frac{5}{2}}_{2,1}(A)$, and thus $u\in C([0,T],B^{\frac{5}{2}}_{2,1}(A))$.
Appendix A Compact Embedding and Time Continuity
Lemma A.1.
Let $s\in\mathbb{R}$. Suppose that $\{u_{N}\}_{N=1}^{\infty}$ is a bounded sequence in the Chemin-Lerner space $\widetilde{L}^{\infty}(0,T;B^{s}_{2,1}(A))$, i.e.,
$\sup_{N\in\mathbb{N}}\sum_{j\in\mathbb{Z}}2^{sj}\|\phi_{j}(D)u_{N}\|_{L^{\infty}(0,T;L^{2})}<\infty.$
Then a subsequence exists such that it converges in the weak-* topology of $\widetilde{L}^{\infty}(0,T;B^{s}_{2,1}(A))$.
Proof. We denote by $c_{0}(\mathbb{Z};L^{1}(0,T;L^{2}))$ the space of sequences $\{f_{j}\}_{j\in\mathbb{Z}}$ of measurable functions on $(0,T)\times\Omega$ such that
$\sup_{j\in\mathbb{Z}}\|f_{j}\|_{L^{1}(0,T;L^{2})}<\infty\quad\text{and}\quad\lim_{|j|\to\infty}\|f_{j}\|_{L^{1}(0,T;L^{2})}=0.$
It is well known that $c_{0}(\mathbb{Z};L^{1}(0,T;L^{2}))$ is separable.
We now consider the mapping
$\widetilde{L}^{\infty}(0,T;B^{s}_{2,1}(A))\ni u\quad\mapsto\quad\{2^{sj}\phi_{j}(D)u\}_{j\in\mathbb{Z}}\in\ell^{1}(\mathbb{Z};L^{\infty}(0,T;L^{2})).$
This map is an isometry, and its range is a closed subspace of $\ell^{1}(\mathbb{Z};L^{\infty}(0,T;L^{2}))$. Note that the space $\widetilde{L}^{\infty}(0,T;B^{s}_{2,1}(A))$ can be identified with a closed subspace of the dual of $\ell^{1}(\mathbb{Z};L^{1}(0,T;L^{2}))$. More precisely, its predual is a quotient space (or closed subspace) related to $c_{0}(\mathbb{Z};L^{1}(0,T;L^{2}))$. For any bounded sequence in $\widetilde{L}^{\infty}(0,T;B^{s}_{2,1}(A))$, the Banach-Alaoglu theorem, combined with the separability of the predual, ensures the existence of a subsequence that converges in the weak-* topology of $\widetilde{L}^{\infty}(0,T;B^{s}_{2,1}(A))$. $\Box$
Lemma A.2.
Let $s\geq 0$. Suppose that $u\in C([0,T],L^{2})\cap\widetilde{L}^{\infty}(0,T;B^{s}_{2,1}(A))$. Then $u\in C([0,T],B^{s}_{2,1}(A))$.
Proof. Let $u$ satisfy the assumptions in Lemma A.2. Suppose $t_{0}\in[0,T]$; we consider the continuity of $u(t)$ at $t=t_{0}$. Let $\varepsilon>0$ be an arbitrary constant. We first choose $J_{0}\in\mathbb{N}$ large enough so that
$\sum_{|j|>J_{0}}2^{sj}\sup_{t\in[0,T]}\|\phi_{j}(D)u(t)\|_{L^{2}}<\varepsilon.$
This choice implies that
$\sum_{|j|>J_{0}}2^{sj}\|\phi_{j}(D)(u(t)-u(t_{0}))\|_{L^{2}}\leq 2\sup_{t\in[0,T]}\sum_{|j|>J_{0}}2^{sj}\|\phi_{j}(D)u(t)\|_{L^{2}}<2\varepsilon$
for all $t\in[0,T]$. On the other hand, the continuity of $u$ in $C([0,T];L^{2})$ and the fact that the sum over $|j|\leq J_{0}$ is a finite sum ensure that there exists $\delta>0$ such that
$\sum_{|j|\leq J_{0}}2^{sj}\|\phi_{j}(D)(u(t)-u(t_{0}))\|_{L^{2}}<\varepsilon$
provided that $|t-t_{0}|<\delta$. Combining these estimates, we conclude that $u(t)\to u(t_{0})$ in $B^{s}_{2,1}(A)$ as $t\to t_{0}$. $\Box$
Acknowledgement
This work was supported by JSPS Grant-in-Aid for Scientific Research (A) JP21H04433.
Declaration of interests The authors declare no conflicts of interest.
Data availability statement Data sharing is not applicable to this article as no data were created or analyzed in this study.